### v2

# Dinkel: Testing Graph Database Engines via State-Aware Query Generation

Abstract.
Graph database management systems (GDBMSs) store and manipulate graph data and form a core part of many data-driven applications. To ensure their reliability, several approaches have been proposed to test GDBMSs by generating queries in Cypher, the most popular graph query language. However, Cypher allows queries with complicated state changes and data dependencies, which existing approaches do not support and thus fail to generate valid, complex queries, thereby missing many bugs in GDBMSs.
In this paper, we propose a novel state-aware testing approach to generate complex Cypher queries for GDBMSs. Our approach models two kinds of graph state, query context and graph schema. Query context describes the available Cypher variables and their corresponding scopes, whereas graph schema summarizes the manipulated graph labels and properties. While generating Cypher queries, we modify the graph states on the fly to ensure each clause within the query can reference the correct state information. In this way, our approach can generate Cypher queries with multiple state changes and complicated data dependencies while retaining high query validity. We implemented this approach as a fully automatic GDBMS testing framework, Dinkel, and evaluated it on three popular open-source GDBMSs, namely Neo4j, RedisGraph, and Apache AGE. In total, Dinkel found 60 bugs, among which 58 were confirmed and 51 fixed. Our evaluation results show that Dinkel can effectively generate complex queries with high validity (93.43%). Compared to existing approaches, Dinkel can cover over 60% more code and find more bugs within the 48-hour testing campaign. We expect Dinkel’s powerful test-case generation to benefit GDBMS testing and help strengthen the reliability of GDBMSs.
copyright: noneccs: Information systems Query languagesccs: Security and privacy Database and storage security
1. Introduction
Graph database management systems (GDBMSs) are crucial for modern interconnected, data-driven computer software. By storing data in graph structures, GDBMSs efficiently manage data through graph properties (e.g., pairs of data connected through certain relationships). Because of their high efficiency, GDBMSs have been widely adopted in large language models (Neo4j and Generative AI, 2023; Unifying LLMs and Knowledge Graphs for GenAI: Use Cases and Best Practices, 2023), recommendation systems (White Paper: Powering Real-Time Recommendations with Graph Database Technology., 2024), social networking (NBC News Analyzes Hundreds of Thousands of Russian Troll Tweets Using Neo4j., 2021), and other data-driven applications (Creating a Modern Solution to Finding What’s Lost with Neo4j., 2024; Novartis Captures the Latest Biological Knowledge for Drug Discovery., 2021). Their practicality has led 75% of the Fortune 100 and all of North America’s top 20 banks to adopt the currently most popular GDBMS, Neo4j (DB-Engines Ranking., 2024; Who Uses Neo4j., 2024).
GDBMSs are rapidly evolving and complex systems (e.g., Neo4j has 1.2M LOC), where bugs are inevitable during their development and maintenance. These bugs are critical as they can corrupt the GDBMSs and lead them to malfunction. The application of error-prone GDBMSs in critical domains may lead to serious consequences if a malicious attack were to abuse a previously unknown GDBMS bug. For example, an attacker can leverage bugs in GDBMSs to disclose confidential data (CVE-2018-1000820., 2018) or run a Denial of Service (DoS) attack by repeatedly crashing the GDBMS server (CVE-2020-35668., 2020).
To improve GDBMS reliability, several testing approaches (Hua et al., 2023; Kamm et al., 2023; Jiang et al., 2023b; Zhuang et al., 2023) have been proposed to find bugs by generating queries in Cypher, the most widely adopted graph query language (openCypher., 2024). As GDBMSs are large-scale, complex systems, effectively testing GDBMSs is difficult and hindered by two key challenges, query generation and test-oracle construction. Specifically, to cover a broader spectrum of functionalities and deeper query-processing logic in GDBMSs, testing approaches need to generate complex queries that involve various Cypher language features and complicated data dependencies. To accurately identify bugs (e.g., logic bugs) triggered by the generated queries, testing approaches need suitable test oracles to validate the correctness of query execution. Query generation and test-oracle construction are orthogonal and equally important for effective GDBMS testing.
However, a significant gap exists in the query generation of existing GDBMS testing approaches. While all approaches focus on integrating test oracles to identify bugs triggered by the generated queries, none systematically models the Cypher query language to improve query generation. As a result, these approaches are limited in generating valid complex queries, i.e., queries involving various clause combinations and complicated data dependencies. Generating such complex queries is challenging because the clauses invoked by Cypher queries can change the graph states of the manipulated graph database. Such state changes are visible in the subsequent clauses. Moreover, the data and variables used by queries have different scopes, depending on the clauses they are involved in. Lacking systematic modeling of these Cypher language features, existing approaches cannot effectively generate complex queries, and thus many critical functionalities and deep logic of GDBMSs are not exercised by the testing campaigns of these approaches. Therefore, many deep bugs (e.g., the bug shown in Figure 3) in GDBMSs are missed by existing approaches.
To generate complex queries for testing GDBMSs, we need to systematically model the Cypher query language. Cypher is a declarative language that allows expressive data querying without describing specific control flows to achieve it, which enables GDBMSs to flexibly select efficient strategies to execute queries. However, different from typical declarative query languages (e.g., SQL), Cypher queries can make changes to the graph database state while executing clauses of the queries, whose effects are visible in subsequent clauses (Cypher Query Language Reference., 2022). For example, a graph node created in an outer CREATE clause can be referenced at other clauses (e.g., DELETE and MERGE) later on in the same query. Such characteristics make it challenging to effectively test GDBMSs. On one hand, neglecting these characteristics tends to make the generated queries simpler, resulting in the GDBMSs not having to handle complicated data dependencies. Therefore, the testing campaign cannot reach the deep logic of GDBMSs. On the other hand, failing to correctly handle such characteristics can make the generated Cypher queries invalid, resulting in many queries being discarded by GDBMSs in early stages (e.g, query parsing). For example, the generated queries become invalid when they reference items that are nonexistent or out-of-scope.
In this paper, we propose, to the best of our knowledge, the first systematic approach to effectively generate complex Cypher queries. Orthogonal to the test-oracle construction (e.g., oracle for detecting logic bugs in GDBMSs), our approach fills the gap in the query generation of GDBMS testing research. The approach models graph states and possible state changes caused by the Cypher queries. It abstracts the states into two categories, query context and graph schemas. Query context contains information related to temporary variables declared in the queries (e.g., the type and scope of each variable), while graph schemas maintain the current graph information, including the graph labels and properties. As query context and graph schemas may change at different Cypher clauses, we must update these abstractions while generating Cypher queries. To this end, we propose state-aware query generation. Instead of determining query skeletons for later expressions complementary, our approach incrementally constructs clauses for the generated queries. When constructing a new clause, our approach references only the elements that are accessible with respect to the current query context and graph schema. After the clause is completed, the approach accordingly updates the state information. In this on-the-fly way, our approach can accurately maintain the dynamically evolving graph state and thus effectively generate queries involving complicated data dependencies and state changes.
Based on our approach, we implemented Dinkel, a fully automatic GDBMS testing framework. As Dinkel is designed to improve query generation, it does not integrate advanced test oracles for finding logic bugs, which is orthogonal to our research. Instead, Dinkel employs only ASan alarms, error messages, and code assertions to identify bugs but already outperforms the state-of-the-art. We have evaluated our tool on three popular GDBMSs (i.e., Neo4j (Neo4j. Graphs for Everyone., 2024), RedisGraph (RedisGraph. A graph database as a Redis module., 2023), and Apache AGE (Apache AGE. Graph database optimized for fast analysis and real-time data processing., 2024)). In the evaluation, Dinkel efficiently generated complex Cypher queries and kept a high validity rate (93.43%). Using these queries, Dinkel found 60 unique, previously unknown bugs. So far, 58 of these bugs have been confirmed, and 51 fixed. Many bugs are long-latent and missed by all existing approaches in their extensive evaluation. Compared to existing approaches, Dinkel can cover over 60% more code and find more bugs within the 48-hour testing campaign, owing to the complex Cypher queries generated by Dinkel. These results demonstrate the effectiveness of Dinkel in generating complex and valid Cypher queries and finding bugs in real-world GDBMSs.
In summary, we make the following technical contributions:
  • Novel approach: We abstract the state information of Cypher queries into query context and graph schema. Based on the abstraction, we propose state-aware query generation, which generates Cypher queries by incrementally constructing Cypher clauses and continuously updating the abstracted states.
  • Practical realization: Based on our approach, we realize a fully automatic and practical testing framework, Dinkel, to find bugs in GDBMSs by generating complex Cypher queries.
  • Promising results: We evaluated Dinkel on three popular open-source GDBMSs, namely Neo4j, RedisGraph, and Apache AGE. Dinkel found 60 unique and previously unknown bugs, among which 58 have been confirmed and 51 have been fixed, demonstrating its effectiveness.
2. Background and Motivation
Table 1. Grammar of Cypher query language
query ::= clause [query]
clause ::= reading_clausewriting_clausereading/writing_clauseprojecting_clause — …
reading_clause ::= ['OPTIONAL'] 'MATCH' pattern ['WHERE' expression]
pattern ::= node [relationship pattern]
node ::= '(' label* properties? ')'
relationship ::= '<-[' label properties? ']-''-[' label properties? ']->'
label ::= ':' identifier
properties ::= '{' identifier : expression [, identifier : expression] '}'
expression ::= identifierconstantoperationfunction — …
writing_clause ::= create_clausedelete_clauseset_clauseremove_clauseforeach_clause — …
reading/writing_clause ::= merge_clausecall_clause — …
projecting_clause ::= return_clausewith_clauseunwind_clause — …
Graph Database Management Systems. Graph database management systems (GDBMSs) utilize graph models to represent the stored data. The most widely used graph model in GDBMSs is the labeled property graph model (Jiang et al., 2023b), which stores interconnected data using nodes connected via relationships (i.e., directed edges between nodes) (openCypher., 2024). Nodes and relationships, so-called graph entities, can be associated with labels and properties. Labels are used to group and classify elements, whereas properties are made up of key-value pairs, providing attribute information of elements.
Figure 1. A property graph modeling a person named Alex owning a car.
Figure 1 shows a simple example represented in a labeled property graph. This example shows a person named Alex owning a vehicle of the type car since 2020. The node on the left-hand side has a label Person and a property Name with a value Alex. This node is connected to another node via a direct edge, i.e. relationship, whose label is OWNS. In this relationship, a property with a key since and a value 2020 is held. The node on the right-hand side has attached the label Vehicle and contains a property named type with value car.
Cypher Query Language. Cypher is the most widely adopted query language for property graph databases (openCypher., 2024) and was proposed by Neo4j, the most popular GDBMS (DB-Engines Ranking., 2024). It is designed as a declarative query language, which allows users to specify the required data without realizing the detailed procedures. The general way to specify data in Cypher is to concretize the graph patterns, which follow the format (n)-[r]->(m) and can be used to reference graph entities that satisfy specified conditions. For example, Figure 2 shows a Cypher query that creates the graph in Figure 1. The query concretizes a graph pattern following the CREATE clause to specify the graph entities to be created.
1CREATE (p:Person {Name: 'Alex'})
2 -[:OWNS {since: 2020}]->
3 (v:Vehicle {type: 'car'});
Figure 2. A Cypher query creating the graph in Figure 1.
Different from another declarative query language, SQL, Cypher does not differentiate between data declaration (DDL), manipulation (DML), and query (DQL) languages. Instead, a Cypher query can create, read, and modify data in a single statement, thereby allowing statements of procedural nature and non-trivial query state manipulation. Generally, Cypher queries access or operate on graphs via clauses, and each query can contain multiple clauses.
Table 1 shows the context-free grammar of Cypher query language. A Cypher query consists of a sequence of clauses. Each clause can be either a reading clause, writing clause, reading/writing clause, projecting clause, or others (e.g., system configuration clause) (Cypher Query Language Reference., 2022). Reading clauses (i.e., MATCH) query the GDBMSs to extract information without modifying the graph entities. Writing clauses (e.g., CREATE, DELETE) modify the data stored in GDBMSs by changing the nodes or relationships in the graph. Reading/writing clauses (e.g., MERGE) can both read and write data in the graph. Projecting clauses (e.g., WITH, RETURN) define expressions to be referenced in the subsequent clauses or the result set. When executing a query with multiple clauses, GDBMSs process these clauses sequentially, and the graph state may change after each clause is processed.
1MERGE (x)<-[:A]-(x)<-[:A]-(x)<-[y:A]->(x)
2DELETE y
3CREATE (x)<-[:B]-()
4DELETE y;
Figure 3. A Cypher query triggering an assertion (”i ¡ n”) failure in RedisGraph.
Figure 3 shows a Cypher query containing multiple clauses. This query triggers an assertion failure in RedisGraph. Figure 4 shows the corresponding graph state changes of Figure 3 after each clause is executed by RedisGraph. The first clause is a MERGE clause that will create graph entities following the specified graph pattern if no graph entity matches the pattern. After this clause is executed, a node x and three relationships with labels A are created. Each relationship is from node x to node x. The subsequent DELETE checks whether the graph entities corresponding to y exist and deletes the existing ones. This clause deletes a relationship. The second CREATE clause creates a new node connected to node x with a new relationship whose label is B. The last clause, DELETE, tries to delete relationship y, but as y has already been deleted and is non-existing, this DELETE clause should do nothing. However, RedisGraph mistakenly recognized y as non-deleted and performed a deletion on a non-existing relationship, which triggered an assertion failure. Such a mistake is caused by the entity ID reuse mechanisms of RedisGraph, which reassigns the ID of the deleted relationship y to the new relationship created by the CREATE clause. When referencing y in the last DELETE clause, RedisGraph mistakenly considers y to still exist because its entity ID is being used.
Cypher queries with multiple clauses changing graph states are commonly used in the real world due to the complex nature of data relationships within graph databases. However, automatically generating such queries (e.g., the query in Figure 3) poses challenges, because (1) the generation needs to be aware of the graph state changes caused by each clause (e.g., MERGE and DELETE); and (2) the intermediate data (e.g., variables x and y) for concretizing these state changes need to be correctly referenced at proper clauses.
xy:A:A:Ax:A:Ax:A:A:BLine 1Line 2Line 3-4
Figure 4. State changes of the query in Figure 3.
Limitation of Existing Approaches. Several approaches have been proposed to test GDBMSs using Cypher (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zhuang et al., 2023; Mang et al., 2024). All of them focus on test oracles to identify bugs triggered by generated queries. They employ only simple template-based query generation, without systematically considering the state changes caused by Cypher clauses. Therefore, their generated queries tend to be structurally simple (i.e., containing fewer clauses) and semantically straightforward (i.e., involving fewer data dependencies). One of the approaches, GDsmith (Hua et al., 2023), integrates techniques to improve the generation of MATCH clauses. It first generates proper graph patterns and WHERE conditions in the MATCH. Then, GDsmith builds a query skeleton according to its grammar template and fills its generated data components (i.e., graph patterns and WHERE conditions). This approach is limited because (1) its query generation is customized for only MATCH clause and cannot be scaled to other clauses; and (2) it still involves no insight for handling the graph state changes caused by Cypher clauses. Therefore, similar to existing approaches, GDsmith is still limited in generating complex queries containing multiple clauses and complicated data dependencies.
Without a systematic model for handling the graph state changes, existing approaches are limited in generating complex Cypher queries. However, many bugs are hidden in the deep logic of DBMS implementations, which may only be covered by complex queries (Jiang et al., 2023a, c). To improve the reliability and security of GDBMSs, proposing an effective approach for generating complex Cypher queries has become imperative and critical for GDBMS testing research.
3. State-Aware GDBMS Testing
In this paper, we propose a novel, state-aware approach for testing GDBMSs using complex Cypher queries. Figure 5 shows the approach overview. Our approach introduces two abstractions, query context and graph schema, to precisely model the graph states maintained by GDBMSs during query processing. Specifically, query context describes the temporary variables declared at each clause along with the corresponding scope and type. Graph schema stores available graph labels and properties at each clause. To generate complex queries that involve multiple clauses, we need to accurately handle the possible state changes caused by these clauses. We propose state-aware query generation, which incrementally constructs Cypher queries clause by clause. Each time a new clause is being constructed, our approach checks the current query context and graph schema, and builds the clause with the available data references. After the clause is constructed, the approach updates the state information accordingly. In this way, our approach can continuously construct Cypher clauses for generating complex queries while accurately maintaining the graph state information for involving complicated data dependencies in the generated queries.
Figure 5. Approach overview.
3.1. Graph State Modeling
Cypher query execution can be affected by two kinds of graph state, which we refer to as query context and graph schema, respectively.
Query Context. One kind of graph state information is the temporary variables declared in the query. These variables have specific types and scopes, affecting only the query where they are declared. They can be either concrete values with specific types (e.g., INTEGER 0) or aliased to specific nodes or relationships (e.g., node x and relationships y in Figure 3). Such variables can be referenced only after they are defined and within their scope. We refer to these variables, their types, and their scope information as query context. Query context may change at different Cypher clauses. Specifically, query context includes new information when a new variable is declared by a clause, and excludes outdated information when the clause is out of the scope of existing variables.
Graph Schema. Another graph state information is the schema of the stored graph data. It includes graph labels and properties (e.g., label A and B for relationships in Figure 3). We refer to such information as graph schema. Graph schemas can be changed by specific Cypher clauses. For example, CREATE clauses can create nodes or relationships with new labels and properties, while REMOVE clauses can remove existing labels or properties.
Different from query context, whose effects are limited by their scope, the operations (e.g., CREATE clause that declares new labels) on graph schema affect the manipulated database permanently. In addition, graph schema can be referenced even though the labels and properties are non-existent in the manipulated graph. This design improves query flexibility as users can write valid queries without concerning the current graph schema.
1CALL {
2 WITH 0 AS x
3 CREATE ()-[:A {n0:x}]->()
4}
5MATCH ()-[:(B & C) {n0: 0}]->(:!(D & E))
6RETURN 0
Figure 6. A Cypher query triggering a Neo4j internal error: ExecutionFailed (Cannot invoke ”RelationshipDataAccessor.type()” because ”this.v3_relationships” is null).
Table 2. State changes of the query in Figure 6
Line Query Context Graph Schema
Label Property
1 {} {} {}
2 {x, INT, CALL} {} {}
3 {x, INT, CALL} {A} {n0}
4 {} {A} {n0}
5 {} {A} {n0}
6 {} {A} {n0}
Example. To better understand query context and graph schema, we discuss the Cypher query shown in Figure 6, which causes an internal error in the enterprise version of Neo4j, which is close-source. This query involves complex graph state changes by using multiple Cypher clauses. Table 2 shows the concrete state at each line of the query. At the first line of the query, query context and graph schema contain no state information as the CALL clause has not had any effects yet. At the second line, a WITH clause declares a temporary variable with value 0, and thus the query context now contains the corresponding variable identifier (i.e., x), its type, and its scope. Because the WITH clause is inside the CALL clause, the scope of the variable x is limited to the CALL clause. At the third line, the CREATE clause updates the graph schema by adding a new relationship with label A and properties n0 referencing variable x. The new graph schema (i.e., label A and property n0) is not limited by the scope of the CALL clause and still exists after the CALL clause and even when the query finishes. At the fifth line, the MATCH clause references the property n0 and four non-existing labels B, C, D, and E. The query ends with a RETURN clause.
Generating such a query is challenging because it involves multiple clauses (5 clauses) and graph data references (e.g, the CREATE references variable x, and the MATCH references property n0). Utilizing the two abstractions, query context and graph schema, our approach can explicitly and precisely describe the graph state maintained by GDBMSs during query processing, which enables our approach to accurately reference available graph data in the generated queries.
1
2
Output :  query
3 Function GenQuery():
4       query $\leftarrow$ EmptyQuery();
       qc $\leftarrow$ {}; // query context
       gs $\leftarrow$ {}; // graph schema
5       do
             // clause can update qc and gs
6             clause, qc, gs $\leftarrow$ GenClause(qc, gs, ANY);
7             AppendQuery(query, clause);
8
9      while rand() ¡ P and Length(query) ¡ Lmax;
10       return query;
11
12Function GenClause(qc, gs, type):
13       if type = ANY then
14             type = RandClauseType();
15
       // initialize the clause according to its type
16       clause, qc, gs $\leftarrow$ RandInitClause(qc, gs, type);
17       foreach subclause, subtype in clause do
18             subclause, qc, gs $\leftarrow$ GenClause(qc, gs, subtype);
19
       // clean out-of-scope query context
20       CleanQC(qc);
21       return clause, qc, gs;
22
Algorithm 1 State-Aware Query Generation
3.2. State-Aware Query Generation
Insight. Our state-aware query generation is based on an observation that both query context and graph schema are updated only when Cypher clauses are invoked (e.g., CREATE and MATCH in Figure 6), or exit specific clause contexts (e.g., CALL in Figure 6). Utilizing this observation, our approach can accurately track the state changes inside the Cypher queries by analyzing the possible impact caused by specific clauses in the queries.
Algorithm. Algorithm 1 shows the procedure of our state-aware query generation. Our approach does not need any input. Initially, it constructs an empty query that contains no clause, and initializes query context qc and graph schema gs to empty sets (line 1-1). Then, the approach incrementally appends the query with a newly generated clause (line 1-1). Each time a new clause is generated, the query context and graph schema may be updated (line 1). Our approach decides whether to append the query with more clauses based on a certain probability P. If the length of the query reaches the limit Lmax, the approach stops appending the query (line 1). In the end, the generated query is returned (line 1).
To generate a clause, our approach needs to reference the query context and graph schema. If the clause type is not specified, the approach randomly chooses a clause type (line 1-1). According to the elements in query context, graph schema, and the specified clause type, our approach randomly initializes a clause (line 1). Initialization for different clauses can vary (e.g., the procedures for CALL and CREATE are different), but it generally includes the procedure of determining the number and types of the components needed by the clause, generating corresponding expressions for the components, and analyzing the impacts of the generated clause. If the generated clause contains subclauses, the approach will recursively call GenClause() to generate subclauses with specified types (line 1-1). After the clause is generated, the approach cleans up the query context if necessary (line 1), such as removing out-of-scope variables at the end of a CALL clause. In the end, the generated clause, the updated query context, and the graph schema are returned (line 1).
Figure 7. Generation process for the query in Figure 6.
Example. Figure 7 shows how our approach generates the query in Figure 6. Initially, the query is empty, and query context and graph schema contain nothing (step ➀). The approach randomly determines that the first clause to be generated is a CALL clause (step ➁). When initializing the CALL clause, our approach determines its components, which are two subclauses. Then, these two subclauses are recursively generated. The first subclause is randomly constructed as a WITH clause with two components, 0 and x. After the WITH subclause is completed, the approach updates the query context by adding a temporary variable x with its type and scope, according to the functionality of WITH (Cypher Query Language Reference., 2022) (step ➂). The approach continues to generate the second subclause, which is randomly determined as a CREATE clause. The component of the CREATE clause is generated with graph pattern ()-[:A {n0:x}]->(), which references the existing variable x. After the CREATE is constructed, the approach updates the graph schema with the newly generated label A and property n0 (step ➃). After the two subclauses are constructed, the CALL generation is completed. Before generating other clauses, our approach cleans the variable x in the query context, because it is scoped to only the CALL clause (step ➄). Then, our approach continues to generate a MATCH clause, which references the existing property n0, according to the latest graph schema, and four non-existing and randomly constructed labels B, C, D, and E (step ➅). In the end, our approach generates a RETURN clause to finish the query (step ➆).
Table 3. Effects of Cypher clauses on graph states
Clause Query Context Graph Schema
Label Property
MATCH + / /
CREATE + + +
MERGE + + +
DELETE / / /
REMOVE / - -
SET / +/- +/-
UNWIND + / /
WITH + / /
RETURN + / /
FOREACH +/- +/- +/-
CALL +/- +/- +/-
UNION +/- +/- +/-
EXISTS +/- +/- +/-
COUNT +/- +/- +/-
  • /: the clause has no effect on the graph state;
    +: the clause can add elements to the graph state;
    -: the clause can remove elements from the graph state;
    +/-: the clause can both add and remove elements.
Clause Impact Analysis. Different Cypher clauses have different effects on query context and graph schema, and our approach models their effects and supports the clause generation accordingly (e.g., RandInitClause function in Algorithm 1). Table 3 lists the effects of different clauses, according to the openCypher specification (Cypher Query Language Reference., 2022). Most clauses can modify either query context, graph schema, or both. The DELETE clause does not affect either query context or graph schemas, because it manipulates only data but does not affect available references or graph attributes. For example, in Figure 3, variable y can still be referenced despite being operated on by a subsequent DELETE clause. Five clauses (i.e., FOREACH, CALL, UNION, EXISTS, and COUNT) can involve subqueries using subclauses, and they can have effects on both query context and graph schemas via invoking specific subclauses. In addition, these clauses demarcate the scope of the query contexts defined by the subclauses. For example, the CALL clause in Figure 6 modifies both query context and the graph schema via the WITH and CREATE clauses, and the variable x defined by the WITH clause can be referenced only within the CALL clause.
4. Implementation
We implement our approach into a practical and fully automatic GDBMS testing framework, Dinkel, integrating Cypher query generation, bug-triggering query reduction, and efficient bug deduplication. Figure 8 shows its architecture. The overall code base of Dinkel consists of 10k lines of Go code. The source code of Dinkel will be available on GitHub. The following discusses important implementation details.
Supported Clauses in Query Generation. As of writing this paper, Dinkel supports all non-administrative clauses (e.g., all the clauses shown in Table 3) in the tested GDBMSs. For example, in Neo4j, 21 out of the total 25 clauses are non-administrative and thereby supported. Administrative clauses, such as SHOW FUNCTIONS and USE, are not supported because they are not related to the logic of query processing in GDBMSs. For example, The clause SHOW FUNCTIONS is designed to list all available functions and cannot be combined with other clauses. The clause USE is used to switch the manipulated databases, which is not necessary as all databases are initialized to empty and they function in the same way.
Compared to existing approaches (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zhuang et al., 2023; Mang et al., 2024), Dinkel supports more Cypher clauses, as shown in Table 4. None of the existing approaches supports FOREACH clauses, because this clause complicates the Cypher control flows and can significantly change graph states during query execution. UNION, EXISTS, and COUNT are not supported as well, as these clauses invoke subqueries that can inherit and change the graph states of the main query. Lacking a systematic model for handling graph state changes, existing approaches are limited in supporting various Cypher clauses.
Bug Detection. In this paper, we focus on the query generation of GDBMS testing and apply only general methods for identifying triggered bugs. Specifically, for each executed query, Dinkel checks the messages returned from the GDBMS. If the message indicates an internal error in the GDBMS (e.g., the error in Figure 6), a bug is identified. We also enable all the assertions embedded in the tested GDBMSs to catch assertion failures. In addition, we enable ASan (Address Sanitizer., 2019) for the GDBMSs developed in C/C++ to identify memory bugs triggered by our generated queries.
Figure 8. Achitecture of Dinkel.
Query Reduction. No existing tool is available for reducing Cypher queries, which makes it difficult for developers to minimize bug-triggering queries and investigate bugs. To address this problem, we implemented Dinkel with an automatic query reduction method. The core idea of this method is to reduce queries clause by clause. For each Cypher clause, Dinkel first tries to delete it. If the query without the clause still triggers the bug, the clause will be permanently removed. Otherwise, the deleted clause will be recovered, and Dinkel goes on to try to replace the clause with an alternative clause if possible. For example, Dinkel can try to replace the CALL clause in Figure 6 with its subclause CREATE. If some clauses are successfully deleted or replaced in one try, Dinkel will restart the reduction process for the reduced queries. The process stops when no clause in the query can be further reduced. With this method, bug-triggering queries can be efficiently reduced.
Bug Deduplication. When a query generated by Dinkel triggers a bug, we need to demonstrate whether this bug is unique or a duplicate of another bug already found by Dinkel. A feasible way is bisection (Git-bisect: Use binary search to find the commit that introduced a bug., 2024; Zeller, 2009), which finds the earliest bug-inducing commits for each triggered bug and deduplicates bugs by checking whether their bug-inducing commits are the same. As GDBMSs are large-scale system software, compiling and building a version of a GDBMS demands considerable time and resources (e.g., 8 minutes for Neo4j in our evaluation), which can make bisection infeasible when a large number of commits need to be bisected (e.g., 79K commits in Neo4j’s GitHub repository (Neo4j. Graphs for Everyone., 2024)). To improve the efficiency of bisections, Dinkel leverages containerization techniques (e.g., Docker (Docker: Accelerated Container Application Development., 2024)) to cache and manage the built GDBMS versions. Given a batch of bugs needed to be deduplicated, Dinkel sets up multiple containers to bisect the bug-inducing commits. During bisection for a bug, Dinkel needs to test many commits by checking whether the bug can be reproduced in this commit version. For each commit that has not been tested yet, Dinkel compiles and builds a GDBMS version, and caches the built version in a container image. When testing a cached commit during bisection for a new bug, Dinkel directly utilizes the image to set up the container, without repeatedly compiling and building the same versions of GDBMSs.
Table 4. Cypher clauses supported by existing approaches
Clause GDsmith GDBMeter GraphGenie GAMERA Dinkel
MATCH
CREATE
MERGE
DELETE
REMOVE
SET
UNWIND
WITH
RETURN
CALL
FOREACH
UNION
EXISTS
COUNT
5. Evaluation
To demonstrate the effectiveness of Dinkel in GDBMS bug detection, we evaluate it on real-world GDBMSs and seek to answer the following questions:
  1. Q1
    Can Dinkel find real bugs in widely-used and extensively-tested GDBMSs? (Section 5.2)
  2. Q2
    How complex and valid are the queries generated by Dinkel? (Section 5.3)
  3. Q3
    How do query context and graph schema contribute to the effectiveness of Dinkel? (Section 5.4)
  4. Q4
    Can Dinkel outperform state-of-the-art GDBMS testing approaches? (Section 5.5)
5.1. Experimental Setup
We evaluated Dinkel on three real-world GDBMSs, Neo4j (Neo4j. Graphs for Everyone., 2024), RedisGraph (RedisGraph. A graph database as a Redis module., 2023), and Apache AGE (Apache AGE. Graph database optimized for fast analysis and real-time data processing., 2024). We chose them because they are popular and open-source GDBMSs supporting Cypher. Neo4j is the most popular GDBMS, according to DB-Engines Ranking (DB-Engines Ranking., 2024), and has been extensively tested by existing approaches (Hua et al., 2023; Jiang et al., 2023b; Kamm et al., 2023; Zhuang et al., 2023). RedisGraph was widely adopted by Redis (the 6th most popular DBMS (DB-Engines Ranking., 2024)) to provide functionalities of graph data querying and storage. Its fork, FalkorDB (FalkorDB. A super fast Graph Database uses GraphBLAS under the hood for its sparse adjacency matrix graph representation., 2024), is also actively developed. Apache AGE is a widely-used extension for PostgreSQL (the 4th most popular DBMS (DB-Engines Ranking., 2024)) that enables applications to manipulate graph data on the top of a relational DBMS.
We evaluate Dinkel on these GDBMSs with the latest versions. During our testing, if the code of the tested GDBMSs is updated (e.g., a new version is released), we set up new Dinkel instances to test the updated versions. Specifically, we test Neo4j starting from version 5.6.0, RedisGraph starting from version 2.12.0, and ApacheAGE starting from version 1.3.0. We clone the code of these GDBMSs from their official GitHub repositories. To demonstrate the effectiveness of Dinkel on testing black-box GDBMSs, we also evaluate Dinkel on the Enterprise version of Neo4j from 5.6.0. Each time we implement new features on Dinkel, we stop and restart the testing. Overall, the testing process lasts three months. The evaluation was performed on Ubuntu 20.04 with a 64-core AMD EPYC 7742 processor running at 2.25GHz and 256GB of RAM.
5.2. Bug Detection
Table 5. Status of the bugs found by Dinkel
GDBMS Reported Confirmed Fixed
Neo4j 32 32 32
RedisGraph 17 17 10
Apache AGE 11 9 9
Total 60 58 51
Table 5 shows the status of the bugs found by Dinkel. In total, Dinkel found 60 unique bugs, including 32 bugs in Neo4j, 17 in RedisGraph, and 11 in Apache AGE. Among these bugs, 58 are confirmed, and 51 are fixed. None of the 60 bugs are duplicates.
Bug Classification. We classify the bugs found by Dinkel into two categories according to their manifestation:
  • Internal errors. The tested GDBMSs unexpectedly throw exceptions or errors when processing syntactically and semantically valid queries. The error messages can indicate the inconsistency of the internal execution status (e.g., the error triggered by the query shown in Figure 6).
  • Crashes. The queries cause GDBMSs to crash due to assertion failures or memory corruption, e.g., a null-pointer dereference.
Table 6. Classifying the found bugs
GDBMS Internal Error Crash
Neo4j 32 0
RedisGraph 3 14
Apache AGE 8 3
Total 43 17
Note that Neo4j is realized in Java and implements exception handling for unexpected errors (e.g., null dereferences in Figure 6 and out-of-bound accesses in Figure 16), and thus Neo4j does not crash but posts exception information when triggering memory corruption. Therefore, we classify all the bugs Dinkel found in Neo4j as internal errors. Table 6 shows the classification results. Among the 60 Dinkel found, 43 cause internal errors, 32 of which are Neo4j bugs. Dinkel can find 17 bugs that crash RedisGraph and Apache AGE, which are implemented in C/C++. Among these 17 bugs, 8 are caused by memory corruptions, and 9 are caused by assertion failures. These results demonstrate that Dinkel is effective in finding bugs in GDBMSs.
Affected GDBMS Components. We investigated 42 of the 51 fixed bugs, where we can analyze the fix patches to identify the GDBMS components affected by the bugs accurately. The 9 fixed bugs in the Enterprise version of Neo4j are not included because developers keep the fix patches confidential. Table 7 shows the results. Among the 42 fixed bugs, 32 affect the parsers of GDBMSs, 15 affect the planners, and 10 affect the executor. Interestingly, 9 bugs can affect two components, and 4 can affect all the components. For example, the bug shown in Figure 16 can affect both the parser and planner of Neo4j. In total, 19 bugs can affect either the planner, the executor, or both. These results demonstrate that (1) Dinkel can find bugs in various GDBMS components; and (2) Dinkel can find bugs in the deep logic of GDBMS implementation, considering 45% (19/42) of bugs are related to the planners or executors.
Table 7. GDBMS components affected by bugs
GDBMS Parser Planner Executor
Neo4j 17 11 6
RedisGraph 7 4 4
Apache AGE 8 0 1
Total 32 15 11
Diversity of Cypher Features. We investigate the queries triggering the 60 bugs and analyze the distribution of the Cypher features used in these queries. All these queries are reduced. The results are shown in Figure 9. The RETURN keyword is present in most bug-triggering queries, appearing in 64% of them. This is unsurprising, as RETURN is required to appear at the end of a query containing only reading clauses (e.g., MATCH), otherwise, the query would be syntactically incorrect. MERGE (33%) and CREATE (26%) are the two clauses inducing the creation of graph properties (e.g., providing available elements for the operations of subsequent clauses). We can deduce from this that graph data is often required to be present for bugs to trigger. The remaining keywords can induce complicated data flow (e.g., WITH (22%) and UNWIND (14%)) or control flow (e.g., CALL (14%) and CASE WHEN (14%)) in the query, indicating that bugs often appear when the database system has to handle such flows. ORDER BY also appears frequently (14%), which forces GDBMSs to keep track of the manipulated graph data and perform non-trivial comparisons to ensure correct order.
Figure 9. Feature distribution of bug-triggering queries.
Data Dependencies for Triggering Bugs. To convey the query complexity Dinkel required to trigger bugs, we analyze the complexity of the 60 bug-triggering queries. The results are shown in Figure 10. Nearly half of the bugs (45%) require at least one data dependency. Specifically, 13 bug-triggering queries contain one data dependency, 8 contain two, and 6 contain three or more. For example, in Figure 3, the bug-triggering query contains 8 data dependencies, as illustrated in Figure 11. Among the 8 dependencies, 6 are involved in query context (i.e., the variables x and y assigned to a node and a relationship), and 2 are involved in graph schema (i.e., the label A). These results demonstrate that some GDBMS bugs can be triggered only when the queries contain multiple data dependencies, and Dinkel can effectively find these bugs.
Figure 10. Data dependencies of bug-triggering queries.
MERGE (x)<-[:A]-(x)<-[:A]-(x)<-[y:A]->(x)DELETEyCREATE (x)<-[:B]-()DELETEy
Figure 11. Data dependencies within the query in Figure 3.
Figure 12. Query sizes of bug-triggering queries.
Size of Bug-Triggering Queries. Figure 12 shows the size of the 60 bug-triggering queries. 54 bugs can be triggered by queries whose size is less than 100 bytes. The bug-triggering queries shown in Figure 3 and Figure 6 are examples of such queries. As query size increases from 0 to 100 bytes, the number of triggered bugs increases almost linearly. The bug-triggering query with the largest size is 146 bytes and is shown in Figure 16. This query triggers an internal error caused by an out-of-bounds access in Neo4j.
Bug Importance. Neo4j, the GDBMS we focus on, provides Community and Enterprise versions. The bugs in the Enterprise version are critical because this version is commonly deployed on commercial applications. Among the 32 Neo4j bugs Dinkel found, 19 bugs can be triggered in both Enterprise and Community versions, and 8 bugs can be triggered in only the Enterprise version. RedisGraph and Apache AGE did not provide severity information about the reported bugs, but we noticed that most of the bugs were fixed, which indicates that the bugs are non-trivial.
Some developers express their appreciation for our effort in finding bugs in their GDBMSs. Particularly, Neo4j provides very positive feedback written in German. We translate it into English:
My colleague has told me that you’ve been busy creating GitHub issues and have contributed to improving Neo4j! Thanks a lot for that! In the name of Neo4j, I would like to send you some swag […]
5.3. Query Generation
To understand the complexity and validity of the queries generated by Dinkel, we ran Dinkel on each tested GDBMS for 24 hours and collected all generated queries. Table 8 shows the statistical results.
Table 8. Queries generated by Dinkel within 24 hours
GDBMS Gen. Valid Clause Dep. Size
Neo4j 68K 89.88% 23.68 30.16 1202.39
RedisGraph 956K 98.14% 23.42 30.15 992.41
Apache AGE 367K 81.79% 16.94 15.60 830.35
Total 1391k 93.43% 21.72 26.31 959.85
  • Gen.: the number of generated queries;
    Valid: the number of valid queries among the generated ones;
    Clause, Dep., and Size: the average number of clauses, data dependencies, and bytes over all generated queries, respectively.
Query Complexity. For the 1391K generated queries, Dinkel constructs 30.23M clauses. The average number of clauses for each query is 21.72. Additionally, on average, each query contains 26.31 data dependencies. Because of the large number of query clauses and data dependencies, the queries generated by Dinkel are typically big. The average query size is 959.85 bytes. These results demonstrate that Dinkel can effectively generate complex queries, which are large and contain multiple clauses with complicated data dependencies. The generated queries are more complex than the bug-triggering queries discussed in Section 5.2, because all the bug-triggering queries are reduced, while the generated queries inevitably contain many redundant parts (Jiang et al., 2023c; Maciver and Donaldson, 2020; Regehr et al., 2012). We find that the queries generated for Apache AGE are less complex (i.e., fewer clauses, fewer data dependencies, and smaller size), because Apache AGE supports fewer clauses, and many clauses that can induce data dependencies are missing. As a result, the relative chance of Dinkel choosing clauses (e.g., RETURN) that terminate query generation becomes much higher, and thus the queries generated for Apache AGE exhibit lower complexity.
Query Validity. Among the 1391K Cypher queries generated by Dinkel, 1300K are valid. The percentage of valid queries is 93.43%. The result demonstrates that Dinkel can keep a high validity rate even when generating complex queries. We investigated the invalid queries generated by Dinkel and found that they were mainly caused by illegal arithmetic operations. Some arithmetic operations require their operands to satisfy some constraints (e.g., the divisor in a division operation must be non-zero). However, Dinkel cannot track the value of each operator in queries because the value may depend on complex expressions, whose results are difficult to calculate (e.g., hash functions). Queries failing to satisfy the constraints of operations cause semantic errors like divisions by zero.
We noticed that the validity of queries generated for Apache AGE is lower than that of the other two GDBMSs. This is caused by the fact that Apache AGE is designed to be a PostgreSQL extension and thus needs to interact with the SQL contexts of PostgreSQL. Specifically, the columns returned by Cypher queries of Apache AGE have to match the type definitions in the wrapping SQL queries used by PostgreSQL, otherwise, the SQL queries become invalid. Figure 13 shows an example, where the Cypher query is valid but the overall SQL query is invalid because the Cypher query returns two columns while the SQL query expects only one column. Such invalid cases do not affect bug detection, because the error of SQL queries occurs only after the Cypher queries are fully executed. For example, in Figure 13, even though the SQL query is invalid, it can crash Apache AGE and PostgreSQL, because the internal Cypher query triggers the bug before it returns values to the SQL query.
1SELECT * FROM cypher('graph',\$\$
2 CREATE (x) SET x.n0 = (true OR true)
3 RETURN 0 AS y, 1 AS z
4\$\$) AS (v agtype);
Figure 13. An invalid SQL query that wraps a valid Cypher query, which crashes Apache AGE and PostgreSQL.
Throughput. On average, one Dinkel instance generates 5.36 test cases per second. Specifically, at each second, one Dinkel instance generates 0.78 test cases for Neo4j, 11 test cases for RedisGraph, and 4.25 test cases for Apache AGE. To understand the bottleneck of test throughput, we further investigated the time used by Neo4j in query generation and query execution, and found that 92.79% of CPU time is occupied by Neo4j executing the generated queries, indicating that the testing throughput is determined by the performance of the tested GDBMS. We believe the current throughput is practical considering (1) GDBMS testing typically lasts for several months (Kamm et al., 2023; Jiang et al., 2023b), and thus a sufficient number of test cases can be executed; and (2) setting up multiple Dinkel instances can significantly improve the test efficiency.
5.4. Ablation Analysis
To understand the contribution of query context and graph schema for bug detection in GDBMSs, we analyze the 60 bug-triggering queries by investigating whether the generation for these queries depends on specific graph states. Specifically, we extract the data dependencies contained by each query and check whether at least one of the data dependencies is related to query context or graph schema. Table 9 shows the analysis results.
Among the 60 bug-triggering queries, 22 depend on only query context, 1 depends on only graph schema, and 4 depend on both (e.g., the bug-triggering query shown in Figure 11). According to these results, without query context, 26 bugs cannot be found. This demonstrates that query context is important because it enables Dinkel to generate queries referencing specific nodes, relationships, or expressions constructed in internal clauses. Without graph schema, a further 5 bugs cannot be found. These bug-triggering queries require Dinkel to properly reference the latest graph labels or properties. The following shows two queries involving query context and graph schema, respectively.
Table 9. Bug-triggering queries related to graph states
Version Found QC GS Both
Neo4j 32 11 1 2
RedisGraph 17 6 0 2
Apache AGE 11 5 0 0
Total 60 22 1 4
  • QC: the queries depend on only query context;
    GS: the queries depend on only graph schema;
    Both: the queries depend on two graph states.
Example 1: Query referencing query context. The query shown in Figure 14 triggers an internal error. This query uses a UNWIND clause to iterate over the elements in the array [0]+[] (i.e., [0]). Then, the query invokes toBoolean to convert the iterated element to the boolean type and return each converted element. However, when concatenating the arrays [0] and [], Neo4j incorrectly maintains the type information of the concatenated array of [0] and []. It directly assigns the type of elements in the right-hand-side array (i.e., []), which is implicit and not determined, to elements of the concatenated array. In the end, the type of the elements in the concatenated array is unexpectedly assigned as Float (while Integer is the expected type), resulting in the internal error of type mismatching. To fix this bug, when concatenating two arrays, developers make Neo4j consider the element type in both arrays.
Example 2: Query referencing graph schema. In Figure 15, the query specifies a graph pattern ()-[:A]->({x:0}), where the relationships with type A are connecting an arbitrary node to another node with property x equal to 0. The MERGE clause tries to find the nodes and relationships satisfying the pattern. If there is no such node and relationship, the MERGE clause creates new ones satisfying the patterns. After the MERGE clause, the EXISTS clause checks whether the graph pattern (:!(B&C))-[{x:0}]->() can be matched in the current graph database and returns the corresponding boolean value. The EXISTS clause references two non-existing labels, B and C, and the property x created in the MERGE clause. As the bug is triggered in the Enterprise version of Neo4j, the root cause and the corresponding fixing patch are kept confidential by Neo4j developers.
1UNWIND [0]+[] AS i
2RETURN toBoolean(i)
Figure 14. A query that references query context and triggers a Neo4j internal error—Type mismatch: expected Boolean, Integer or String but was Float (line 1, column 37 (offset: 36))
1MERGE ()-[:A]->({x:0})
2RETURN EXISTS {(:!(B&C))-[{x:0}]->()};
Figure 15. A query that references graph schema and triggers a Neo4j internal error—Neo4jError: ExecutionFailed (Cannot invoke ”RelationshipDataAccessor.properties(…)” because ”relationshipCursor” is null).
5.5. Comparison
Bug Latency Study. To demonstrate that Dinkel can find bugs missed by existing approaches (Kamm et al., 2023; Hua et al., 2023; Jiang et al., 2023b; Zhuang et al., 2023; Mang et al., 2024), we investigate the latencies of the bugs found by Dinkel. For each bug, we check whether its bug-inducing commit was created before the years when existing approaches were published. If Dinkel finds some long-latent bugs, we can conclude that Dinkel can find bugs missed by existing approaches. This comparison is reasonable and objective because: (1) none of the bugs found by Dinkel are marked as duplicated by GDBMS developers, which means that no approach found these bugs until Dinkel found them; and (2) all existing approaches have extensively tested Neo4j and RedisGraph (Kamm et al., 2023; Hua et al., 2023; Jiang et al., 2023b; Zhuang et al., 2023; Mang et al., 2024), which means that in these two GDBMSs, no approach found the long-latent bugs found by Dinkel during their evaluation.
GDsmith (Hua et al., 2023), GDBMeter (Kamm et al., 2023), GraphGenie (Jiang et al., 2023b), and GAMERA (Zhuang et al., 2023) were published in 2023, and GRev (Mang et al., 2024) was published in 2024. Therefore, for each bug, we checked whether its bug-inducing commit was created before 2024 (i.e., in 2023 or earlier) and 2023. Table 10 shows the results. Among the 49 bugs Dinkel found in Neo4j and RedisGraph, 33 bugs already existed before 2023, indicating that all the existing approaches proposed after 2023 failed to find these 33 bugs in their extensive evaluation. In addition, 14 bugs were introduced to Neo4j and RedisGraph between 2023 and 2024, while GRev, which was proposed after 2024, cannot find these bugs. These results indicate that existing approaches indeed miss many long-latent bugs, while Dinkel can effectively find them.
Table 10. Latency of the logic bugs found by Dinkel
DBMS Found Bug-involved year Longest latency
¡ 2024 ¡ 2023
Neo4j 32 30 20 2016
RedisGraph 17 17 13 2020
Apache AGE 11 11 7 2021
Total 60 58 40 2016
We analyze the bug-triggering queries of these bugs and conclude the reason why existing approaches missed these bugs as the following: (1) the bug-triggering queries (e.g., the queries shown in Figure 3 and Figure 6) contain data dependencies, which existing approaches lack support for handling; and (2) the bug-triggering queries contain Cypher features (e.g., CALL and FOREACH in the query shown in Figure 16) that are not supported by existing approaches. The following is an example of the 33 bugs.
1WITH [] AS n0 ORDER BY null
2CALL {
3 WITH [] AS n1 ORDER BY null
4 UNWIND [0] AS x
5 UNWIND [x] AS n2
6 RETURN n2 AS n3
7}
8FOREACH (n4 IN null | MERGE ({key:n3}))
Figure 16. A query triggering a Neo4j internal error—Neo4j-
Error: ExecutionFailed (arraycopy: last destination index 6 out of bounds for object array[5]).
Example: Query using data dependencies and advanced Cypher features. Figure 16 shows a bug in Neo4j, which is present in 2022 but none of the existing approaches found it. The bug-triggering query involves complicated data dependencies and advanced Cypher clauses (e.g., FOREACH and UNWIND). This query constructs an empty array [] and invokes a subquery using the CALL clause. The subquery also constructs an empty array. It then invokes two UNWIND clauses, which iterate over each element in the operated array. For each iterated element, the UNWIND executes the subsequent clause under the context of this element. For example, the array [x] used by the second UNWIND references the variable x, which is 0 when the first UNWIND iterates over the item 0 in the array [0]. After the CALL, the query utilizes FOREACH, whose execution depends on the query context of the subquery in the CALL. To optimize the query execution, Neo4j tries to flatten the FOREACH loop. However, such optimization does not work well when the loop involves update operations (i.e., MERGE) under complicated contexts (i.e., the contexts produced by CALL). The improper optimization corrupts the internal data structures of Neo4j, Eagers, which are the production of another optimization, Eagerness analysis. In the end, an internal error is triggered when Neo4j tries to access the corrupted Eagers. To fix this bug, Neo4j developers modify both the Eagerness analysis and the flattening strategy for FOREACH clauses to ensure they work consistently.
Empirical Comparison. To empirically demonstrate the effectiveness of Dinkel on code coverage and bug detection over the state-of-the-art, we evaluated Dinkel and existing Cypher testing tools (i.e., GDSmith, GDBMeter, GraphGenie, GAMERA, and GRev) on Neo4j and RedisGraph, which are the only two GDBMSs supported by all these tools. Each evaluation persisted for 48 hours. Table 11 shows the comparison results of Dinkel and existing tools.
Table 11. Results of comparison
Neo4j RedisGraph
Covered Line Bug Covered Line Bug
GDsmith 533k 0 15k 0
GDBMeter 283k 1 19k 1
GraphGenie 506k 0 3k 0
GAMERA 283k 1 3k 0
GRev 601k 1 14k 1
Dinkel 737k 11 20k 18
Figure 17. Covered lines of Neo4j and RedisGrpah.
Code Coverage. As shown in Table 11, on average, Dinkel can cover 67% and 85% more code than existing approaches in Neo4j and RedisGraph, respectively. Among the tools except for Dinkel, GRev covers the most code (i.e., 601k lines) in Neo4j. Compared to GRev, Dinkel covers 23% more code. Benefiting from the powerful capability of generating complex Cypher queries, Dinkel can cover much deeper logic of GDBMSs that is related to processing advanced Cypher features and complicated data dependencies. In contrast, simple queries generated by existing tools have little chance to touch such logic. In RedisGraph, except for Dinkel, GDBMeter covers the most code (i.e., 19k lines). Compared to GDBMeter, Dinkel covers 5% more code. The coverage improvement is less significant than in Neo4j, because RedisGraph supports fewer Cypher features and cannot handle some complicated Cypher semantics. For example, RedisGraph does not support processing the subqueries in FOREACH and CALL clauses. As a result, the space for Dinkel to improve the code coverage in RedisGraph is limited. Figure 17 shows the trends of the covered code in each GDBMS during testing. Dinkel and all the state-of-the-art cover new codes quickly and saturate in earlier stages. In almost all testing campaigns, Dinkel can cover more code than the existing tools.
Found Bugs. As shown in Table 11, compared to existing approaches, Dinkel finds more bugs in both Neo4j and RedisGraph. We describe the relations of the found bugs in Figure 18. Note that none of these approaches have found logic bugs within 48 hours, so the bugs shown are related to internal errors or crashes. Specifically, Dinkel found 11 bugs in Neo4j. GDBMeter, GRev, and GAMERA found 1 bug, and this bug can also be found by Dinkel. In RedisGraph, Dinkel found 17 bugs, while GRev and GDBMeter found only 1 bug, and this bug can also be found by Dinkel. For the 27 bugs missed by existing approaches, their bug-triggering queries contain either complicated data dependencies or advanced Cypher features that are not supported by existing approaches. The bug shown in Figure 16 is one of the 27 bugs. These results demonstrate that Dinkel can find more bugs than existing approaches by generating more complex Cypher queries.
Figure 18. Relation of bugs found by existing tools.
6. Related Work
GDBMS Testing. GDBMS testing is an emerging research field and several approaches (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zheng et al., 2022; Zhuang et al., 2023; Mang et al., 2024) are proposed. All these approaches integrate techniques for constructing test oracles to identify bugs, especially logic bugs (i.e., the bugs cause GDBMSs to return incorrect results). Grand (Zheng et al., 2022) targets GDBMSs using the Gremlin query language and detects logic bugs using differential testing (McKeeman, 1998). Several approaches (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zhuang et al., 2023; Mang et al., 2024) support GDBMSs using Cypher. Similar to Grand, GDsmith (Hua et al., 2023) also uses differential testing to find logic bugs. GDBMeter (Kamm et al., 2023) leverages the idea of query partitioning (Rigger and Su, 2020b), which decomposes the predicate of a query and checks whether the queries with the decomposed predicates produce consistent results with the original query. GraphGenie (Jiang et al., 2023b) modifies the graph patterns used in a query and checks whether the query with modified graph patterns satisfies the expected relationship (i.e., equivalence, subset, or superset) with the original queries. GAMERA (Zhuang et al., 2023) extracts metamorphic relations (Chen et al., 1998) from the manipulated graph data and validates GDBMSs by checking whether the outputs of generated queries satisfy these relations.
Different from all these existing approaches, which focus on new test oracles for finding logic bugs, Dinkel focuses on improving query generation, which is the groundwork of GDBMS testing. An interesting future work is to combine the techniques of Dinkel and existing approaches, enhancing both query generation and test-oracle construction for finding more bugs in GDBMSs.
RDBMS Testing. Compared to GDBMS testing, testing relational database management systems (RDBMSs) is more mature, where both query generation (SQLsmith: A random SQL query generator., 2023; Zhong et al., 2020; Fu et al., 2022; Jiang et al., 2023a; Liang et al., 2023) and test-oracle construction (Rigger and Su, 2020c, a, b; Hao et al., 2023; Song et al., 2023; Tang et al., 2023; Jiang and Su, 2024) are well-researched. We focus on discussing the work on query generation for RDBMSs. SQLsmith (SQLsmith: A random SQL query generator., 2023) embeds the SQL grammar (SQL standard., 1992) and can generate queries containing complex statements. SQUIRREL (Zhong et al., 2020) integrates a new intermediate representation (IR) for modeling the dependencies among statements in a SQL query. Therefore, SQUIRREL can generate queries containing multiple statements. DynSQL (Jiang et al., 2023a) incrementally generates SQL statements for a query. For each statement, DynSQL queries the tested RDBMS to capture the latest DBMS state. In this way, DynSQL can generate complex and valid queries with multiple statements. LEGO (Liang et al., 2023) proposes type-affinity to describe the pattern of composing two types of SQL statements. Based on the type-affinity extracted from existing queries, LEGO can mutate and synthesize new queries that have a higher chance of increasing the code coverage of RDBMSs.
Different from these approaches, Dinkel is designed for GDBMSs and integrates techniques for improving Cypher query generation.
State-Aware Fuzzing. Fuzzing is a promising technique for detecting bugs in various software (Jiang et al., 2020, 2022; American Fuzzy Lop., 2021; Lyu et al., 2019; Yun et al., 2018; Padhye et al., 2019; Chen and Chen, 2018). Some approaches (Ma et al., 2023; Kim et al., 2022; Jiang et al., 2023a; Zhao et al., 2022; Atlidakis et al., 2019) have been proposed to find bugs more efficiently in state-sensitive systems. RESTler (Atlidakis et al., 2019) analyzes the API specification of the tested cloud service and generates request sequences that follow the inferred producer-consumer dependencies. RESTler also collects the response observed during prior executed requests to guide subsequent request sequence generation. LOKI (Ma et al., 2023) proposed to test blockchain consensus protocols. It builds a state model to dynamically track the state transition of each node in the blockchain systems and accordingly generates inputs with proper targets, types, and contents. StateFuzz (Zhao et al., 2022) is designed to test Linux drivers. It utilizes static analysis to recognize critical variables that affect the control flows or memory accesses, and represents program states using these variables. StateFuzz prioritizes test cases triggering new states using three-dimensional feedback. To effectively test USB gadget stacks, FuzzUSB (Kim et al., 2022) extracts the internal state machines from USB gadget drivers via static analysis and symbolic execution, and uses such state information as fuzzing feedback to guide test-case generation.
Similar to these approaches, Dinkel also leverages state information and is the first state-aware testing approach designed for detecting bugs in GDBMSs.
7. Conclusion
In this paper, we have presented a novel, practical approach for testing GDBMSs using complex Cypher queries. Our approach models the graph state into two categories: query context and graph schema. Based on the model, we have proposed state-aware query generation, which incrementally constructs the clauses for the generated queries and intermittently updates the state information after each clause is constructed. We have implemented our approach as an automatic GDBMS testing framework, Dinkel, and evaluated it on three popular GDBMSs. The evaluation results demonstrate that Dinkel can effectively generate complex Cypher queries with high validity. Dinkel found 60 bugs, many of which had been missed by existing approaches. Considering its effectiveness in Cypher query generation, we believe Dinkel can be the foundation for GDBMS testing, facilitating and inspiring follow-up research, like Csmith (Yang et al., 2011) for compiler testing.

### v3

@@
-Dinkel: Testing Graph Database Engines via State-Aware Query Generation
+Dinkel: State-Aware and Granular Framework for Validating Graph Databases
 Abstract.
-Graph database management systems (GDBMSs) store and manipulate graph data and form a core part of many data-driven applications.
-To ensure their reliability, several approaches have been proposed to test GDBMSs by generating queries in Cypher, the most popular graph query language.
-However, Cypher allows queries with complicated state changes and data dependencies, which existing approaches do not support and thus fail to generate valid, complex queries, thereby missing many bugs in GDBMSs.
-In this paper, we propose a novel state-aware testing approach to generate complex Cypher queries for GDBMSs.
-Our approach models two kinds of graph state, query context and graph schema.
-Query context describes the available Cypher variables and their corresponding scopes, whereas graph schema summarizes the manipulated graph labels and properties.
-While generating Cypher queries, we modify the graph states on the fly to ensure each clause within the query can reference the correct state information.
-In this way, our approach can generate Cypher queries with multiple state changes and complicated data dependencies while retaining high query validity.
-We implemented this approach as a fully automatic GDBMS testing framework, Dinkel, and evaluated it on three popular open-source GDBMSs, namely Neo4j, RedisGraph, and Apache AGE.
-In total, Dinkel found 60 bugs, among which 58 were confirmed and 51 fixed.
-Our evaluation results show that Dinkel can effectively generate complex queries with high validity (93.43%).
-Compared to existing approaches, Dinkel can cover over 60% more code and find more bugs within the 48-hour testing campaign.
-We expect Dinkel’s powerful test-case generation to benefit GDBMS testing and help strengthen the reliability of GDBMSs.
+Graph database management systems (GDBMSs) have been powering many data-driven applications.
+To ensure GDBMS reliability, several testing approaches have been proposed.
+However, they all suffer from two key limitations: (1) insufficient support for generating complex and valid queries to exercise deep GDBMS code, and (2) lack of general oracles to validate the execution correctness of arbitrary queries.
+In this paper, we propose a novel and practical approach Dinkel, to thoroughly test GDBMSs.
+Our approach consists of two core techniques.
+First, to generate complex and valid queries, we model two kinds of graph state, query context and graph schema, to describe the Cypher variables and the manipulated graph labels and properties.
+We generate queries clause-by-clause, and modify the graph states on the fly to ensure each clause references the correct state information.
+Second, to generally validate query results, we introduce two fine-grained query transformations: clause-level and expression-level transformations.
+These transformations can operate on arbitrary queries while preserving their semantics.
+Dinkel validates GDBMSs by checking whether the transformed query produces the same results as the original.
+We evaluated Dinkel on three well-known GDBMSs.
+In total, we found 127 bugs, among which 113 were confirmed, 84 were fixed, and 33 were logic bugs.
+Compared to existing approaches, Dinkel can cover over 70% more code and find substantially more bugs within a 48-hour testing campaign.
+We expect Dinkel’s powerful bug detection to lay a practical foundation for GDBMS testing.
 copyright: noneccs: Information systems Query languagesccs: Security and privacy Database and storage security
 1. Introduction
-Graph database management systems (GDBMSs) are crucial for modern interconnected, data-driven computer software.
-By storing data in graph structures, GDBMSs efficiently manage data through graph properties (e.g., pairs of data connected through certain relationships).
-Because of their high efficiency, GDBMSs have been widely adopted in large language models (Neo4j and Generative AI, 2023; Unifying LLMs and Knowledge Graphs for GenAI: Use Cases and Best Practices, 2023), recommendation systems (White Paper: Powering Real-Time Recommendations with Graph Database Technology., 2024), social networking (NBC News Analyzes Hundreds of Thousands of Russian Troll Tweets Using Neo4j., 2021), and other data-driven applications (Creating a Modern Solution to Finding What’s Lost with Neo4j., 2024; Novartis Captures the Latest Biological Knowledge for Drug Discovery., 2021).
-Their practicality has led 75% of the Fortune 100 and all of North America’s top 20 banks to adopt the currently most popular GDBMS, Neo4j (DB-Engines Ranking., 2024; Who Uses Neo4j., 2024).
-GDBMSs are rapidly evolving and complex systems (e.g., Neo4j has 1.2M LOC), where bugs are inevitable during their development and maintenance.
-These bugs are critical as they can corrupt the GDBMSs and lead them to malfunction.
-The application of error-prone GDBMSs in critical domains may lead to serious consequences if a malicious attack were to abuse a previously unknown GDBMS bug.
-For example, an attacker can leverage bugs in GDBMSs to disclose confidential data (CVE-2018-1000820., 2018) or run a Denial of Service (DoS) attack by repeatedly crashing the GDBMS server (CVE-2020-35668., 2020).
-To improve GDBMS reliability, several testing approaches (Hua et al., 2023; Kamm et al., 2023; Jiang et al., 2023b; Zhuang et al., 2023) have been proposed to find bugs by generating queries in Cypher, the most widely adopted graph query language (openCypher., 2024).
-As GDBMSs are large-scale, complex systems, effectively testing GDBMSs is difficult and hindered by two key challenges, query generation and test-oracle construction.
-Specifically, to cover a broader spectrum of functionalities and deeper query-processing logic in GDBMSs, testing approaches need to generate complex queries that involve various Cypher language features and complicated data dependencies.
-To accurately identify bugs (e.g., logic bugs) triggered by the generated queries, testing approaches need suitable test oracles to validate the correctness of query execution.
-Query generation and test-oracle construction are orthogonal and equally important for effective GDBMS testing.
-However, a significant gap exists in the query generation of existing GDBMS testing approaches.
-While all approaches focus on integrating test oracles to identify bugs triggered by the generated queries, none systematically models the Cypher query language to improve query generation.
-As a result, these approaches are limited in generating valid complex queries, i.e., queries involving various clause combinations and complicated data dependencies.
-Generating such complex queries is challenging because the clauses invoked by Cypher queries can change the graph states of the manipulated graph database.
-Such state changes are visible in the subsequent clauses.
-Moreover, the data and variables used by queries have different scopes, depending on the clauses they are involved in.
-Lacking systematic modeling of these Cypher language features, existing approaches cannot effectively generate complex queries, and thus many critical functionalities and deep logic of GDBMSs are not exercised by the testing campaigns of these approaches.
-Therefore, many deep bugs (e.g., the bug shown in Figure 3) in GDBMSs are missed by existing approaches.
-To generate complex queries for testing GDBMSs, we need to systematically model the Cypher query language.
-Cypher is a declarative language that allows expressive data querying without describing specific control flows to achieve it, which enables GDBMSs to flexibly select efficient strategies to execute queries.
-However, different from typical declarative query languages (e.g., SQL), Cypher queries can make changes to the graph database state while executing clauses of the queries, whose effects are visible in subsequent clauses (Cypher Query Language Reference., 2022).
-For example, a graph node created in an outer CREATE clause can be referenced at other clauses (e.g., DELETE and MERGE) later on in the same query.
+Graph database management systems (GDBMSs) are crucial for modern interconnected, data-driven computer software (Creating a Modern Solution to Finding What’s Lost with Neo4j., 2024; Novartis Captures the Latest Biological Knowledge for Drug Discovery., 2021; Neo4j and Generative AI, 2023; NBC News Analyzes Hundreds of Thousands of Russian Troll Tweets Using Neo4j., 2021; Yin et al., 2023; Chen and Qian, 2023; Xu et al., 2019).
+Notably, 75% of the Fortune 100 and all of North America’s top 20 banks have adopted the currently most popular GDBMS, Neo4j (DB-Engines Ranking., 2024; Who Uses Neo4j., 2024).
+While rapidly evolving and becoming more complex (e.g., Neo4j has 1.2M LOC), GDBMSs inevitably suffer from bugs introduced during their development.
+These bugs are critical as they can corrupt the GDBMSs and lead to serious consequences.
+One of the most critical types of bugs are logic bugs, which can silently cause GDBMSs to produce incorrect query results.
+To improve GDBMS reliability, testing approaches (Hua et al., 2023; Kamm et al., 2023; Jiang et al., 2023b; Zhuang et al., 2023) have been proposed to find bugs by generating queries in Cypher, the most widely-adopted graph query language (openCypher., 2024).
+As large-scale, complex systems, GDBMSs are difficult to test effectively due to two key challenges: query generation and test-oracle construction.
+To cover a broad spectrum of functionalities and deep system logic, approaches need to generate complex queries involving various Cypher language features and complicated data dependencies.
+To identify bugs, especially logic bugs that cause GDBMSs to silently produce incorrect results for queries, testing approaches require test oracles that can be generally applied to validate arbitrary queries.
+However, both of these challenges remain unsolved.
+All existing approaches focus on proposing new test oracles, but struggle to systematically generate complex and valid queries, leaving much GDBMS logic and code inadequately tested or completely unexercised.
+Although existing approaches focus on test oracles to identify logic bugs, none of the proposed oracles can generally be applied to validate arbitrary queries.
+As a result, many logic bugs in GDBMSs get missed, even though they can be triggered by existing queries.
+The following discusses the details of these two challenges.
+Challenge 1: Generating Complex Queries.
+In DBMSs, many bugs can only be triggered by complex queries (Jiang et al., 2023a; Jiang and Su, 2024; Jiang et al., 2023c).
+A complex Cypher query can involve various clause combinations and complicated data dependencies.
+Generating such complex queries is challenging because the clauses invoked by Cypher queries can change the graph states of the manipulated graph database during their execution.
+Such state changes are visible in subsequent clauses.
+Moreover, the data and variables used by Cypher queries have different scopes, depending on the clauses they are involved in.
+Due to the lack of systematic modeling of these Cypher language features, existing approaches cannot effectively generate complex queries, and thus, many critical functionalities and deep logic of GDBMSs are not exercised by the testing campaigns of these approaches.
+Therefore, many deep bugs (e.g., the bugs shown in Figure 1 and Figure 2) in GDBMSs are missed by existing approaches.
+Modeling Cypher language features is not trivial.
+Cypher is a declarative language that allows expressive data querying without describing specific control flows, which enables GDBMSs to flexibly select efficient strategies to execute queries.
+However, different from typical declarative query languages (e.g., SQL), Cypher queries can make changes to the graph database state while executing Cypher clauses of the queries, whose effects are visible in subsequent clauses (Cypher Query Language Reference., 2022).
+For example, a graph node created in an outer CREATE clause can be referenced by subsequent DELETE clauses.
 Such characteristics make it challenging to effectively test GDBMSs.
-On one hand, neglecting these characteristics tends to make the generated queries simpler, resulting in the GDBMSs not having to handle complicated data dependencies.
-Therefore, the testing campaign cannot reach the deep logic of GDBMSs.
-On the other hand, failing to correctly handle such characteristics can make the generated Cypher queries invalid, resulting in many queries being discarded by GDBMSs in early stages (e.g, query parsing).
-For example, the generated queries become invalid when they reference items that are nonexistent or out-of-scope.
-In this paper, we propose, to the best of our knowledge, the first systematic approach to effectively generate complex Cypher queries.
-Orthogonal to the test-oracle construction (e.g., oracle for detecting logic bugs in GDBMSs), our approach fills the gap in the query generation of GDBMS testing research.
-The approach models graph states and possible state changes caused by the Cypher queries.
+On the one hand, neglecting these characteristics tends to make the generated queries simpler, resulting in the GDBMSs not having to handle complicated data dependencies and therefore not allowing the testing campaign to reach the deep logic of GDBMSs.
+On the other hand, failing to correctly handle such characteristics can make the generated Cypher queries invalid, resulting in many queries being discarded by GDBMSs in the early stages (e.g., query parsing).
+For example, the generated queries may become invalid when referencing nonexistent or out-of-scope items.
+MERGE (x)<-[:A]-(x)<-[:A]-(x)<-[y:A]->(x)
+DELETE y
+CREATE (x)<-[:B]-()
+DELETE y;
+Figure 1. Assertion failure found by Dinkel in RedisGraph.
+Challenge 2: General Test Oracle.
+Existing approaches use query transformations to validate the correctness of query results (Kamm et al., 2023; Jiang et al., 2023b; Zhuang et al., 2023; Mang et al., 2024a).
+Given a generated Cypher query, they transform it into another query and check whether the results of the transformed query and the original query follow specified relationships.
+However, these approaches cannot be generally applied to arbitrary queries because they require their generated queries to follow specific query patterns so that their transformations can be applied.
+For example, GraphGenie (Jiang et al., 2023b) requires the predicate in MATCH clauses to satisfy the preconditions of their transformation rules (e.g., one rule requires the predicates to involve a cyclic graph).
+Moreover, GraphGenie can transform and validate only queries that contain a single MATCH clause followed by a RETURN.
+GRev (Mang et al., 2024a) integrates an abstract syntax graph (ASG) for interpreting the predicate in MATCH statements and can thus transform the predicate in more diverse ways and with fewer restrictions.
+However, the predicate modeling in GRev can only be applied to MATCH statements, while queries involving other common clauses like MERGE, UNWIND, and FOREACH remain unsupported and thus untested.
+Such inapplicabilities arise from the coarse-grained transformation used by existing approaches.
+Specifically, their transformations are at the query level.
+When transforming a query to another related query, these approaches need to take into account the semantics of the entire query and choose transformation rules according to these parsed semantics.
+They cannot handle semantics that do not fit their transformation rules (e.g., both GraphGenie and GRev do not support FOREACH clauses).
+Such coarse-grained transformations make existing approaches inapplicable to various queries that do not match specific query patterns.
+Therefore, many logic bugs are missed.
+For example, existing approaches miss the bug shown in Figure 2 because the original query is too complex and does not follow any specified patterns.
+// Original: 2 rows
+RETURN 0 AS n1
+UNION CALL {
+FOREACH (n2 IN [] |
+FOREACH (n3 IN [] |
+MERGE (:A {u:0})))
+} MATCH (y)
+RETURN y AS n1
+UNION CREATE ()
+RETURN 0 AS n1
+// Transformed: 1 row
+FOREACH (n0 IN [] |
+FOREACH (x IN [] |
+SET x = {}))
+RETURN 0 AS n1
+UNION CALL {
+FOREACH (n2 IN [] |
+FOREACH (n3 IN [] |
+MERGE (:A {u:0})))
+} MATCH (y)
+RETURN y AS n1
+UNION CREATE ()
+RETURN 0 AS n1
+Figure 2. Queries exposing a logic bug in Neo4j.
+In this paper, we propose a novel and practical approach, Dinkel, which is designed to lay a practical foundation for testing GDBMSs.
+Dinkel integrates two technical solutions to address the two challenges discussed above.
+Solution 1: State-Aware Query Generation.Dinkel models graph states and state changes caused by Cypher queries.
 It abstracts the states into two categories, query context and graph schemas.
@@
 To this end, we propose state-aware query generation.
-Instead of determining query skeletons for later expressions complementary, our approach incrementally constructs clauses for the generated queries.
-When constructing a new clause, our approach references only the elements that are accessible with respect to the current query context and graph schema.
-After the clause is completed, the approach accordingly updates the state information.
-In this on-the-fly way, our approach can accurately maintain the dynamically evolving graph state and thus effectively generate queries involving complicated data dependencies and state changes.
-Based on our approach, we implemented Dinkel, a fully automatic GDBMS testing framework.
-As Dinkel is designed to improve query generation, it does not integrate advanced test oracles for finding logic bugs, which is orthogonal to our research.
-Instead, Dinkel employs only ASan alarms, error messages, and code assertions to identify bugs but already outperforms the state-of-the-art.
-We have evaluated our tool on three popular GDBMSs (i.e., Neo4j (Neo4j.
-Graphs for Everyone., 2024), RedisGraph (RedisGraph.
-A graph database as a Redis module., 2023), and Apache AGE (Apache AGE.
-Graph database optimized for fast analysis and real-time data processing., 2024)).
-In the evaluation, Dinkel efficiently generated complex Cypher queries and kept a high validity rate (93.43%).
-Using these queries, Dinkel found 60 unique, previously unknown bugs.
-So far, 58 of these bugs have been confirmed, and 51 fixed.
-Many bugs are long-latent and missed by all existing approaches in their extensive evaluation.
-Compared to existing approaches, Dinkel can cover over 60% more code and find more bugs within the 48-hour testing campaign, owing to the complex Cypher queries generated by Dinkel.
-These results demonstrate the effectiveness of Dinkel in generating complex and valid Cypher queries and finding bugs in real-world GDBMSs.
-In summary, we make the following technical contributions:
-  • Novel approach: We abstract the state information of Cypher queries into query context and graph schema.
-Based on the abstraction, we propose state-aware query generation, which generates Cypher queries by incrementally constructing Cypher clauses and continuously updating the abstracted states.
-  • Practical realization: Based on our approach, we realize a fully automatic and practical testing framework, Dinkel, to find bugs in GDBMSs by generating complex Cypher queries.
-  • Promising results: We evaluated Dinkel on three popular open-source GDBMSs, namely Neo4j, RedisGraph, and Apache AGE.
-Dinkel found 60 unique and previously unknown bugs, among which 58 have been confirmed and 51 have been fixed, demonstrating its effectiveness.
-2. Background and Motivation
+Instead of determining query skeletons for later expressions complementarily, our approach incrementally constructs clauses for the generated queries.
+When constructing a new Cypher clause, our approach references only the accessible elements according to the current query context and graph schema.
+After the clause is completed, the approach updates the corresponding state information.
+In this on-the-fly way, our approach can accurately maintain the dynamically evolving graph state and thus generate queries involving complicated data dependencies and state changes.
+Solution 2: Fine-Grained Query Transformation.
+Our idea to validate Cypher queries is inspired by EET (Jiang and Su, 2024), which is designed for SQL queries in relational DBMSs (RDBMSs).
+EET proposes to validate SQL queries by transforming them at the level of expressions.
+In this paper, we further generalize this method to fine-grained query transformation, without limiting the transformation units to just expressions.
+Specifically, instead of analyzing the semantics of a whole query, our approach transforms queries by their fine-grained units.
+In the context of Cypher queries, the fine-grained units can not only be expressions, but also clauses, which can flexibly change graph states.
+By operating on these low-level units, we focus on the fine-grained semantics of queries, without the need to analyze the overall query-level semantics.
+For example, we can transform the first RETURN clause of the original query in Figure 2 by inserting an additional FOREACH clause, which is dead code and has no effects because the array to be iterated over is empty.
+Therefore, the clauses before and after transformations should be semantically equivalent, and the transformed query should produce the same results as the original one.
+However, their results differ, indicating that a logic bug was triggered.
+In this process, Dinkel only needs to ensure the semantic equivalence between clauses, without understanding the semantics of the whole query.
 Table 1. Grammar of Cypher query language
@@
 projecting_clause ::= return_clausewith_clauseunwind_clause — …
+Regarding the transformation for clauses, Dinkel manipulates target clauses to affect the graph state in ways that do not interfere with the semantics of the following clauses (e.g., creating and immediately deleting a node), or introduce dead code (e.g., a MATCH clause matching a node that does not exist) before target clauses.
+Regarding the transformation for expressions, given an arbitrary expression within clauses, Dinkel can construct a semantically equivalent expression by leveraging logical (e.g., De Morgan’s laws) and algebraic equivalences (e.g., $x\equiv x+0$).
+Dinkel also uses features specific to the Cypher language for transforming expressions (e.g., x $\equiv$ CASE WHEN true THEN x END).
+Ultimately, we implemented Dinkel as a fully automatic GDBMS testing framework.
+Dinkel will be open-sourced once the paper is ready to be public.
+We have evaluated our tool on three popular GDBMSs (i.e., Neo4j (Neo4j.
+Graphs for Everyone., 2024), RedisGraph (RedisGraph.
+A graph database as a Redis module., 2023), and Memgraph (Memgraph.
+Open-source graph database, tuned for dynamic analytics environments.
+Easy to adopt, scale and own., 2024)).
+In the evaluation, Dinkel efficiently generated complex Cypher queries and kept a high validity rate (89.02%).
+Using these queries, Dinkel found 127 unique, previously unknown bugs, among which 33 are logic bugs.
+So far, 113 of these bugs have been confirmed, and 84 fixed.
+Many bugs are long-latent and missed by all existing approaches during their extensive evaluation.
+Compared to existing approaches, Dinkel can cover over 70% more code and find more bugs within the 48-hour testing campaign.
+These results demonstrate that Dinkel significantly outperforms existing approaches in bug detection for GDBMSs.
+In summary, we make the following contributions:
+  • Novel approach: We tackle two fundamental problems of GDBMS testing via two novel solutions: (1) state-aware query generation, which abstracts the Cypher state information to facilitate query generation; and (2) fine-grained query transformation, which can be generally applied to validate arbitrary Cypher queries by manipulating their clauses and expressions.
+  • Practical realization: We realize a fully automatic testing framework, Dinkel, to find bugs in GDBMSs by generating and validating complex Cypher queries.
+  • Promising results: We evaluated Dinkel on three popular open-source GDBMSs, namely Neo4j, RedisGraph and Memgraph.
+Dinkel found 127 unique and new bugs, among which 33 are logic bugs.
+So far, 113 unique bugs have been confirmed, and 84 fixed.
+These results demonstrate the effectiveness of Dinkel.
+2. Background
 Graph Database Management Systems.
-Graph database management systems (GDBMSs) utilize graph models to represent the stored data.
-The most widely used graph model in GDBMSs is the labeled property graph model (Jiang et al., 2023b), which stores interconnected data using nodes connected via relationships (i.e., directed edges between nodes) (openCypher., 2024).
-Nodes and relationships, so-called graph entities, can be associated with labels and properties.
-Labels are used to group and classify elements, whereas properties are made up of key-value pairs, providing attribute information of elements.
-Figure 1. A property graph modeling a person named Alex owning a car.
-Figure 1 shows a simple example represented in a labeled property graph.
-This example shows a person named Alex owning a vehicle of the type car since 2020.
-The node on the left-hand side has a label Person and a property Name with a value Alex.
-This node is connected to another node via a direct edge, i.e.
-relationship, whose label is OWNS.
-In this relationship, a property with a key since and a value 2020 is held.
-The node on the right-hand side has attached the label Vehicle and contains a property named type with value car.
-Cypher Query Language.
-Cypher is the most widely adopted query language for property graph databases (openCypher., 2024) and was proposed by Neo4j, the most popular GDBMS (DB-Engines Ranking., 2024).
+GDBMSs utilize graph models to represent data.
+The most widely used graph model is the labeled property graph model (Jiang et al., 2023b), which stores interconnected data using nodes connected via relationships (i.e., directed edges between nodes) (openCypher., 2024).
+Nodes and relationships, generally referred to as graph entities, can be associated with labels and properties.
+Labels are used to group and classify elements, whereas properties are made up of key-value pairs, providing attribute information.
+Cypher Language.
+Cypher is the most widely adopted query language for property graph databases (openCypher., 2024) and was proposed by Neo4j.
 It is designed as a declarative query language, which allows users to specify the required data without realizing the detailed procedures.
-The general way to specify data in Cypher is to concretize the graph patterns, which follow the format (n)-[r]->(m) and can be used to reference graph entities that satisfy specified conditions.
-For example, Figure 2 shows a Cypher query that creates the graph in Figure 1.
-The query concretizes a graph pattern following the CREATE clause to specify the graph entities to be created.
-1CREATE (p:Person {Name: 'Alex'})
-2 -[:OWNS {since: 2020}]->
-3 (v:Vehicle {type: 'car'});
-Figure 2. A Cypher query creating the graph in Figure 1.
-Different from another declarative query language, SQL, Cypher does not differentiate between data declaration (DDL), manipulation (DML), and query (DQL) languages.
-Instead, a Cypher query can create, read, and modify data in a single statement, thereby allowing statements of procedural nature and non-trivial query state manipulation.
+The general way to specify data in Cypher is to concretize the graph patterns, which follow the format (n)-[r]->(m) and can be used to reference graph entities satisfying specified conditions.
+For example, the query in Figure 1 concretizes a graph pattern following the CREATE clause to specify the graph entities to be created.
+Unlike the query language SQL, Cypher does not differentiate between data declaration (DDL), manipulation (DML), or query (DQL) languages.
+Instead, a Cypher query can create, read, and modify data in a single query, thereby allowing queries of procedural nature and non-trivial state manipulation.
 Generally, Cypher queries access or operate on graphs via clauses, and each query can contain multiple clauses.
-Table 1 shows the context-free grammar of Cypher query language.
+Table 1 shows the context-free grammar of the Cypher query language.
 A Cypher query consists of a sequence of clauses.
 Each clause can be either a reading clause, writing clause, reading/writing clause, projecting clause, or others (e.g., system configuration clause) (Cypher Query Language Reference., 2022).
-Reading clauses (i.e., MATCH) query the GDBMSs to extract information without modifying the graph entities.
+Reading clauses (i.e., MATCH) query the GDBMS to extract information without modifying the graph entities.
 Writing clauses (e.g., CREATE, DELETE) modify the data stored in GDBMSs by changing the nodes or relationships in the graph.
@@
 When executing a query with multiple clauses, GDBMSs process these clauses sequentially, and the graph state may change after each clause is processed.
-1MERGE (x)<-[:A]-(x)<-[:A]-(x)<-[y:A]->(x)
-2DELETE y
-3CREATE (x)<-[:B]-()
-4DELETE y;
-Figure 3. A Cypher query triggering an assertion (”i ¡ n”) failure in RedisGraph.
-Figure 3 shows a Cypher query containing multiple clauses.
+xy:A:A:Ax:A:Ax:A:A:BLine 1Line 2Line 3-4
+Figure 3. State changes of the query in Figure 1.
+Illustrative Example.
+Figure 1 shows a query with multiple clauses.
 This query triggers an assertion failure in RedisGraph.
-Figure 4 shows the corresponding graph state changes of Figure 3 after each clause is executed by RedisGraph.
-The first clause is a MERGE clause that will create graph entities following the specified graph pattern if no graph entity matches the pattern.
-After this clause is executed, a node x and three relationships with labels A are created.
+Figure 3 shows the corresponding graph state changes during execution of each clause.
+The first clause, MERGE, will create graph entities following the specified graph pattern if no graph entity matches the pattern.
+Therefore, a node x and three relationships with labels A are created.
 Each relationship is from node x to node x.
-The subsequent DELETE checks whether the graph entities corresponding to y exist and deletes the existing ones.
-This clause deletes a relationship.
-The second CREATE clause creates a new node connected to node x with a new relationship whose label is B.
-The last clause, DELETE, tries to delete relationship y, but as y has already been deleted and is non-existing, this DELETE clause should do nothing.
-However, RedisGraph mistakenly recognized y as non-deleted and performed a deletion on a non-existing relationship, which triggered an assertion failure.
+The subsequent DELETE checks whether the graph entity y exists and deletes the existing ones, ultimately deleting one relationship.
+The second CREATE clause creates a new node connected to node x with a new relationship B.
+The last clause, DELETE, tries to delete relationship y, but as y has already been deleted, this DELETE should do nothing.
+However, RedisGraph recognized y as non-deleted and performed a deletion on a nonexistent relationship, which triggered an assertion failure.
 Such a mistake is caused by the entity ID reuse mechanisms of RedisGraph, which reassigns the ID of the deleted relationship y to the new relationship created by the CREATE clause.
@@
 Cypher queries with multiple clauses changing graph states are commonly used in the real world due to the complex nature of data relationships within graph databases.
-However, automatically generating such queries (e.g., the query in Figure 3) poses challenges, because (1) the generation needs to be aware of the graph state changes caused by each clause (e.g., MERGE and DELETE); and (2) the intermediate data (e.g., variables x and y) for concretizing these state changes need to be correctly referenced at proper clauses.
-xy:A:A:Ax:A:Ax:A:A:BLine 1Line 2Line 3-4
-Figure 4. State changes of the query in Figure 3.
-Limitation of Existing Approaches.
-Several approaches have been proposed to test GDBMSs using Cypher (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zhuang et al., 2023; Mang et al., 2024).
-All of them focus on test oracles to identify bugs triggered by generated queries.
-They employ only simple template-based query generation, without systematically considering the state changes caused by Cypher clauses.
-Therefore, their generated queries tend to be structurally simple (i.e., containing fewer clauses) and semantically straightforward (i.e., involving fewer data dependencies).
-One of the approaches, GDsmith (Hua et al., 2023), integrates techniques to improve the generation of MATCH clauses.
-It first generates proper graph patterns and WHERE conditions in the MATCH.
-Then, GDsmith builds a query skeleton according to its grammar template and fills its generated data components (i.e., graph patterns and WHERE conditions).
-This approach is limited because (1) its query generation is customized for only MATCH clause and cannot be scaled to other clauses; and (2) it still involves no insight for handling the graph state changes caused by Cypher clauses.
-Therefore, similar to existing approaches, GDsmith is still limited in generating complex queries containing multiple clauses and complicated data dependencies.
-Without a systematic model for handling the graph state changes, existing approaches are limited in generating complex Cypher queries.
-However, many bugs are hidden in the deep logic of DBMS implementations, which may only be covered by complex queries (Jiang et al., 2023a, c).
-To improve the reliability and security of GDBMSs, proposing an effective approach for generating complex Cypher queries has become imperative and critical for GDBMS testing research.
-3. State-Aware GDBMS Testing
-In this paper, we propose a novel, state-aware approach for testing GDBMSs using complex Cypher queries.
-Figure 5 shows the approach overview.
-Our approach introduces two abstractions, query context and graph schema, to precisely model the graph states maintained by GDBMSs during query processing.
-Specifically, query context describes the temporary variables declared at each clause along with the corresponding scope and type.
-Graph schema stores available graph labels and properties at each clause.
-To generate complex queries that involve multiple clauses, we need to accurately handle the possible state changes caused by these clauses.
-We propose state-aware query generation, which incrementally constructs Cypher queries clause by clause.
-Each time a new clause is being constructed, our approach checks the current query context and graph schema, and builds the clause with the available data references.
-After the clause is constructed, the approach updates the state information accordingly.
-In this way, our approach can continuously construct Cypher clauses for generating complex queries while accurately maintaining the graph state information for involving complicated data dependencies in the generated queries.
+However, it is extremely challenging to automatically generate and validate such queries, because (1) the generation needs to be aware of the graph state changes caused by each clause (e.g., MERGE and DELETE) and properly reference the intermediate data (e.g., variables x and y); and (2) the validation needs to construct general test oracles in the case of the complicated semantics the query can contain.
+3. Approach
+We propose Dinkel, a GDBMS testing framework integrating two techniques: state-aware query generation and fine-grained query transformation. Figure 4 shows its overview.
+3.1. State-Aware Query Generation
+Dinkelintroduces two abstractions, query context and graph schema, to precisely model the graph states maintained by GDBMSs.
+Based on these abstractions, Dinkel incrementally constructs Cypher queries clause by clause.
+Each time a new clause is being constructed, our approach checks the current query context and graph schema, builds the clause with the available data references, and updates the state information accordingly.
+In this way, Dinkel continuously constructs clauses for generating complex queries while accurately maintaining the graph state information to introduce complicated data dependencies.
 Figure 4. Approach overview.
-3.1. Graph State Modeling
-Cypher query execution can be affected by two kinds of graph state, which we refer to as query context and graph schema, respectively.
+3.1.1. Graph State Modeling.
+Cypher query execution can be affected by two kinds of graph states:
 Query Context.
-One kind of graph state information is the temporary variables declared in the query.
-These variables have specific types and scopes, affecting only the query where they are declared.
-They can be either concrete values with specific types (e.g., INTEGER 0) or aliased to specific nodes or relationships (e.g., node x and relationships y in Figure 3).
-Such variables can be referenced only after they are defined and within their scope.
+One kind of graph state are the temporary variables declared in the query.
+These variables have specific types and scopes, affecting only the query state where they are declared.
+They can either be concrete values with specific types (e.g., integer 0) or aliased to specific nodes or relationships (e.g., node x and relationship y in Figure 1).
+Such variables can be referenced only after they are defined, within their scope.
 We refer to these variables, their types, and their scope information as query context.
@@
 Graph Schema.
-Another graph state information is the schema of the stored graph data.
+Another graph state is the schema of the stored graph data.
 It includes graph labels and properties (e.g., label A and B for relationships in Figure 1).
-We refer to such information as graph schema.
-Graph schemas can be changed by specific Cypher clauses.
+We refer to such state information as graph schema.
+Graph schemas can be changed by specific clauses.
 For example, CREATE clauses can create nodes or relationships with new labels and properties, while REMOVE clauses can remove existing labels or properties.
-Different from query context, whose effects are limited by their scope, the operations (e.g., CREATE clause that declares new labels) on graph schema affect the manipulated database permanently.
-In addition, graph schema can be referenced even though the labels and properties are non-existent in the manipulated graph.
-This design improves query flexibility as users can write valid queries without concerning the current graph schema.
-1CALL {
-2 WITH 0 AS x
-3 CREATE ()-[:A {n0:x}]->()
-4}
-5MATCH ()-[:(B & C) {n0: 0}]->(:!(D & E))
-6RETURN 0
-Figure 6. A Cypher query triggering a Neo4j internal error: ExecutionFailed (Cannot invoke ”RelationshipDataAccessor.type()” because ”this.v3_relationships” is null).
-Table 2. State changes of the query in Figure 6
-Line Query Context Graph Schema
-Label Property
-1 {} {} {}
-2 {x, INT, CALL} {} {}
-3 {x, INT, CALL} {A} {n0}
-4 {} {A} {n0}
-5 {} {A} {n0}
-6 {} {A} {n0}
-Example.
-To better understand query context and graph schema, we discuss the Cypher query shown in Figure 6, which causes an internal error in the enterprise version of Neo4j, which is close-source.
-This query involves complex graph state changes by using multiple Cypher clauses.
-Table 2 shows the concrete state at each line of the query.
-At the first line of the query, query context and graph schema contain no state information as the CALL clause has not had any effects yet.
-At the second line, a WITH clause declares a temporary variable with value 0, and thus the query context now contains the corresponding variable identifier (i.e., x), its type, and its scope.
-Because the WITH clause is inside the CALL clause, the scope of the variable x is limited to the CALL clause.
-At the third line, the CREATE clause updates the graph schema by adding a new relationship with label A and properties n0 referencing variable x.
-The new graph schema (i.e., label A and property n0) is not limited by the scope of the CALL clause and still exists after the CALL clause and even when the query finishes.
-At the fifth line, the MATCH clause references the property n0 and four non-existing labels B, C, D, and E.
-The query ends with a RETURN clause.
-Generating such a query is challenging because it involves multiple clauses (5 clauses) and graph data references (e.g, the CREATE references variable x, and the MATCH references property n0).
-Utilizing the two abstractions, query context and graph schema, our approach can explicitly and precisely describe the graph state maintained by GDBMSs during query processing, which enables our approach to accurately reference available graph data in the generated queries.
+Different from query context, whose effects are limited by their scope, the operations (e.g., CREATE clause that declares new labels) on graph schema affect the database permanently.
+In addition, graph schema can be referenced even though the labels and properties are nonexistent.
+This design improves query flexibility as users can write valid queries without concerning themselves with the current graph schema.
+3.1.2. Query Generation
+Our generation is based on a key insight that clauses are the basic units for operating on graphs, and both query context and graph schema are updated only when clauses are invoked or exit specific clause context.
+Based on this insight, our approach track the state changes inside Cypher queries by analyzing the possible impact caused by specific clauses in queries.
 1
@@
 5 do
-             // clause can update qc and gs
 6    clause, qc, gs $\leftarrow$ GenClause(qc, gs, ANY);
-7             AppendQuery(query, clause);
+7    AppendClause(query, clause);
 8
@@
 15
-       // initialize the clause according to its type
+ // initialize the clause accordingly
 16 clause, qc, gs $\leftarrow$ RandInitClause(qc, gs, type);
-17       foreach subclause, subtype in clause do
-18             subclause, qc, gs $\leftarrow$ GenClause(qc, gs, subtype);
+17 foreach subclause, subclause_type in clause do
+18    subclause, qc, gs $\leftarrow$ GenClause(qc, gs, subclause_type);
 19
+20
  // clean out-of-scope query context
@@
 22 return clause, qc, gs;
-22
+23
 Algorithm 1 State-Aware Query Generation
-3.2. State-Aware Query Generation
-Insight.
-Our state-aware query generation is based on an observation that both query context and graph schema are updated only when Cypher clauses are invoked (e.g., CREATE and MATCH in Figure 6), or exit specific clause contexts (e.g., CALL in Figure 6).
-Utilizing this observation, our approach can accurately track the state changes inside the Cypher queries by analyzing the possible impact caused by specific clauses in the queries.
 Algorithm.
@@
 Our approach decides whether to append the query with more clauses based on a certain probability P.
-If the length of the query reaches the limit Lmax, the approach stops appending the query (line 1).
+If the length of the query reaches the limit Lmax, the approach stops appending to the query (line 1).
 In the end, the generated query is returned (line 1).
-To generate a clause, our approach needs to reference the query context and graph schema.
+To generate a clause, our approach references the query context and graph schema.
 If the clause type is not specified, the approach randomly chooses a clause type (line 1-1).
 According to the elements in query context, graph schema, and the specified clause type, our approach randomly initializes a clause (line 1).
-Initialization for different clauses can vary (e.g., the procedures for CALL and CREATE are different), but it generally includes the procedure of determining the number and types of the components needed by the clause, generating corresponding expressions for the components, and analyzing the impacts of the generated clause.
+Initialization for different clauses can vary (e.g., the procedures for CALL and CREATE are different), but it generally includes the procedure of determining the number of components needed by the clause and their types, generating corresponding expressions for these components, and analyzing the impacts of the generated clause.
 If the generated clause contains subclauses, the approach will recursively call GenClause() to generate subclauses with specified types (line 1-1).
@@
 In the end, the generated clause, the updated query context, and the graph schema are returned (line 1).
-Figure 7. Generation process for the query in Figure 6.
-Example.
-Figure 7 shows how our approach generates the query in Figure 6.
-Initially, the query is empty, and query context and graph schema contain nothing (step ➀).
-The approach randomly determines that the first clause to be generated is a CALL clause (step ➁).
-When initializing the CALL clause, our approach determines its components, which are two subclauses.
-Then, these two subclauses are recursively generated.
-The first subclause is randomly constructed as a WITH clause with two components, 0 and x.
-After the WITH subclause is completed, the approach updates the query context by adding a temporary variable x with its type and scope, according to the functionality of WITH (Cypher Query Language Reference., 2022) (step ➂).
-The approach continues to generate the second subclause, which is randomly determined as a CREATE clause.
-The component of the CREATE clause is generated with graph pattern ()-[:A {n0:x}]->(), which references the existing variable x.
-After the CREATE is constructed, the approach updates the graph schema with the newly generated label A and property n0 (step ➃).
-After the two subclauses are constructed, the CALL generation is completed.
-Before generating other clauses, our approach cleans the variable x in the query context, because it is scoped to only the CALL clause (step ➄).
-Then, our approach continues to generate a MATCH clause, which references the existing property n0, according to the latest graph schema, and four non-existing and randomly constructed labels B, C, D, and E (step ➅).
-In the end, our approach generates a RETURN clause to finish the query (step ➆).
-Table 3. Effects of Cypher clauses on graph states
-Clause Query Context Graph Schema
-Label Property
-MATCH + / /
-CREATE + + +
-MERGE + + +
-DELETE / / /
-REMOVE / - -
-SET / +/- +/-
-UNWIND + / /
-WITH + / /
-RETURN + / /
-FOREACH +/- +/- +/-
-CALL +/- +/- +/-
-UNION +/- +/- +/-
-EXISTS +/- +/- +/-
-COUNT +/- +/- +/-
-  • /: the clause has no effect on the graph state;
-    +: the clause can add elements to the graph state;
-    -: the clause can remove elements from the graph state;
-    +/-: the clause can both add and remove elements.
-Clause Impact Analysis.
-Different Cypher clauses have different effects on query context and graph schema, and our approach models their effects and supports the clause generation accordingly (e.g., RandInitClause function in Algorithm 1).
-Table 3 lists the effects of different clauses, according to the openCypher specification (Cypher Query Language Reference., 2022).
-Most clauses can modify either query context, graph schema, or both.
-The DELETE clause does not affect either query context or graph schemas, because it manipulates only data but does not affect available references or graph attributes.
-For example, in Figure 3, variable y can still be referenced despite being operated on by a subsequent DELETE clause.
-Five clauses (i.e., FOREACH, CALL, UNION, EXISTS, and COUNT) can involve subqueries using subclauses, and they can have effects on both query context and graph schemas via invoking specific subclauses.
-In addition, these clauses demarcate the scope of the query contexts defined by the subclauses.
-For example, the CALL clause in Figure 6 modifies both query context and the graph schema via the WITH and CREATE clauses, and the variable x defined by the WITH clause can be referenced only within the CALL clause.
+Figure 5. Generation for the original query in Figure 2.
+3.1.3. Illustrative Example
+We show the effectiveness of our generation using a complex query example, the original query in Figure 2.
+As shown in Figure 5, initially, the query is empty, and both query context and graph schema contain nothing (step ➀).
+The approach randomly determines that the first clause is a RETURN clause with an integer 0 as the returned value n1.
+Then, the approach updates the query context by adding n1 with its type and scope (step ➁).
+The second clause is a UNION clause.
+After constructing the UNION, the approach cleans the query context as the clause after the UNION cannot reference the variables defined before the UNION (step ➂).
+Subsequently, the approach randomly decides to generate a CALL clause with a subclause (step ➃).
+The subclause is recursively generated as a FOREACH clause, defining n2 in its loop condition.
+The FOREACH also has a subclause, which is another FOREACH clause, defining n3.
+The last FOREACH has a MERGE as its subclause.
+Our approach records the query context (i.e., n2, n3 defined by the two FOREACH clauses) and graph schema (i.e., label A and property u referenced by the MERGE) updated in this recursive generation process (step ➄).
+After completing generation of the FOREACH clauses, the approach removes n2 and n3 from query context as they are scoped to only the FOREACH clauses.
+Finishing the CALL, our approach continues to randomly generate clauses, including a MATCH defining y, and a RETURN referencing y as n1, and update the query context by adding y and n1 (step ➅).
+Finally, our approach randomly generates a UNION followed by a CREATE and a RETURN.
+The UNION cleans all the contents of the query context, after which RETURN adds a new variable n1 (step ➆).
+3.2. Fine-Grained Query Transformation
+To enable general query validation, we construct test oracles based on the fine-grained units of queries: clauses and expressions.
+Each query consists of a chain of clauses, which can be transformed by modifying the clauses or prepending them with new clauses.
+Each clause may include expressions, on which more transformations can be applied by using algebraic identities or Cypher features.
+Table 2.
+Clause-level transformation rules.
+(1) T.S.: transformation strategies including Dead Code (DC), Unused Context (UC), Redundant Write (RW), and Inconsequential Supersession (IS); (2) <literal>: any atomic expression; (3) <mirror(x)>: a mirrored version of x (e.g., a->b $\Rightarrow$ b<-a); (4) <clauses>: a random number of random clauses; (5) <new pattern>: a graph pattern that involves at least one previously unused label or property
+Type Original T.S. Transformed
+Read- Only C DC CALL {MATCH x WHERE false <clauses>} C
+DC CALL {MATCH <new pattern> <clauses>} C
+IS OPTIONAL MATCH <new pattern> C
+UC CALL {WITH * RETURN x AS i} C
+UC UNWIND <literal> AS i C
+WITH x UC WITH x, y AS i
+MATCH x IS MATCH x WHERE true
+IS MATCH <mirror(x)>
+Write/ Project C DC FOREACH (i IN [] |<clauses>) C
+RW DELETE <deleted var> C
+RW CREATE x DELETE x C
+RW SET x += {} C
+REMOVE x RW REMOVE x, x
+DELETE x RW DELETE x, x
+IS DETACH DELETE x
+Any C IS WITH * C
+3.2.1. Clause-Level Transformations
+These transformations manipulate queries by either adding new clauses, or switching out clauses with equivalent counterparts.
+Table 3.2 lists the transformation rules.
+We design different rules for reading clauses and writing/projecting clauses because the Cypher specification limits the kinds of clauses that can be used before specific clauses.
+The transformations consider state information to avoid semantic invalidity (e.g., repeatedly defining the same variables).
+We introduce four strategies to achieve semantic-preserving transformation for clauses: Dead Code (DC), Unused Context (UC), Redundant Write (RW), and Inconsequential Supersession (IS).
+Dead Code expands the target clause by introducing additional clauses that, due to Cypher’s way of handling queries, will not be executed.
+We implement three rules for this strategy.
+Two rules transform read clauses by adding a CALL clause with a MATCH.
+The MATCH clause is designed to return no rows, leading to no rows for the clauses following the MATCH to ingest.
+As a result, the subsequent clauses in the CALL clauses will not be executed.
+One rule transforms write clauses by inserting a FOREACH before them.
+The FOREACH iterates the elements in a given array and invokes clauses to operate on the iterated elements.
+Because the array is empty, no iterations are performed by the FOREACH.
+Therefore, the invoked clauses in the FOREACH will not be executed.
+For example, in Figure 2, our approach transforms the first RETURN clause of the original query by adding a dead FOREACH, which preserves the semantics of the RETURN and thus should not affect the final results.
+However, the results of the transformed query change, indicating that a bug was triggered.
+Unused Context expands a clause by involving unused query context.
+Specifically, it defines new variables that are not referenced by subsequent parts of the query.
+Redundant Write is straightforward.
+It expands a clause by deleting a previously deleted node, removing the same property multiple times, or creating and immediately deleting a subgraph.
+Lastly, Inconsequential Supersession expands a clause by adding inconsequential operations (e.g., adding WHERE true to a MATCH clause) or modifying the formats of the clauses (e.g., reversing the graph pattern of a MATCH clause).
+Clause-level transformations have the ability to influence the query context through manipulating variables, as well as changing the database state by modifying graph elements or affecting control flow.
+By doing so, Dinkel can effectively force the GDBMSs to execute system logic that was not covered by the original queries, and thus validate the query results by checking whether the different executed logic produces the same results.
+Table 3.
+Expression-level transformation rules.
+(1) T.S.: transformation strategies including Mathematical Identity (MI) and Cypher Feature (CF); (2) P': randomly generated boolean expression; (3) E': randomly generated expression
+Type Initial T.S. Transformed
+Bool P MI NOT (NOT P)
+MI P OR false
+MI P AND true
+P OR Q MI NOT ((NOT P) AND (NOT Q))
+P AND Q MI NOT ((NOT P) OR (NOT Q))
+true MI true OR P'
+MI P' OR (NOT P') OR (P' IS NULL)
+false MI false AND P'
+MI P' AND (NOT P') AND (P' IS NOT NULL)
+NULL MI NULL {XOR|=|<>|<|>|>=|<=} P'
+Num x MI -(-x)
+MI x {+|-} 0
+MI x {*|/|^} 1
+0 MI 0 * x
+NULL MI NULL {+|-|*|/} x
+Str S MI S + ””
+List L MI L + []
+Any E IS NULL MI NOT (E IS NOT NULL)
+E CF CASE WHEN true THEN E ELSE E' END
+CF CASE WHEN false THEN E' ELSE E END
+3.2.2. Expression-Level Transformations
+These transformations operate on expressions within Cypher clauses.
+Although they do not modify the state of the database or query, they still force the GDBMS to evaluate more complex expressions.
+Table 3.2.1 lists our expression-level transformations.
+Our expression-level transformations involve two strategies: Mathematical Identity (MI) and Cypher Feature (CF).
+Mathematical Identity leverages mathematical laws (e.g., De Morgan’s laws for boolean con- and disjunctions, and double negations) to transform the expressions while preserving their semantics.
+Cypher Feature leverages the features of the Cypher language.
+For example, Dinkel can transform an arbitrary expression into a CASE expression, where the branch condition is determined to be true, and the corresponding value of the true branch is the original expression.
+When transforming expressions, Dinkel sometimes needs to involve new components in the transformed expressions (e.g., the E' in the transformed CASE expressions).
+In this case, Dinkel constructs these components by referencing the available variables recorded in the corresponding query context.
 4. Implementation
-We implement our approach into a practical and fully automatic GDBMS testing framework, Dinkel, integrating Cypher query generation, bug-triggering query reduction, and efficient bug deduplication.
-Figure 8 shows its architecture.
+We implement our approach as a fully automatic GDBMS testing framework, Dinkel.
 The overall codebase of Dinkel consists of 10k lines of Go code.
-The source code of Dinkel will be available on GitHub.
+The source code of Dinkel is available for review on an anonymized repository [^1] https://github.com/Anon10214/dinkel and will be public concurrent with the publication of the paper.
 The following discusses important implementation details.
 Supported Clauses in Query Generation.
-As of writing this paper, Dinkel supports all non-administrative clauses (e.g., all the clauses shown in Table 3) in the tested GDBMSs.
+Dinkel supports all non-administrative clauses in the tested GDBMSs.
 For example, in Neo4j, 21 out of the total 25 clauses are non-administrative and thereby supported.
-Administrative clauses, such as SHOW FUNCTIONS and USE, are not supported because they are not related to the logic of query processing in GDBMSs.
-For example, The clause SHOW FUNCTIONS is designed to list all available functions and cannot be combined with other clauses.
-The clause USE is used to switch the manipulated databases, which is not necessary as all databases are initialized to empty and they function in the same way.
-Compared to existing approaches (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zhuang et al., 2023; Mang et al., 2024), Dinkel supports more Cypher clauses, as shown in Table 4.
-None of the existing approaches supports FOREACH clauses, because this clause complicates the Cypher control flows and can significantly change graph states during query execution.
-UNION, EXISTS, and COUNT are not supported as well, as these clauses invoke subqueries that can inherit and change the graph states of the main query.
-Lacking a systematic model for handling graph state changes, existing approaches are limited in supporting various Cypher clauses.
-Bug Detection.
-In this paper, we focus on the query generation of GDBMS testing and apply only general methods for identifying triggered bugs.
-Specifically, for each executed query, Dinkel checks the messages returned from the GDBMS.
-If the message indicates an internal error in the GDBMS (e.g., the error in Figure 6), a bug is identified.
-We also enable all the assertions embedded in the tested GDBMSs to catch assertion failures.
-In addition, we enable ASan (Address Sanitizer., 2019) for the GDBMSs developed in C/C++ to identify memory bugs triggered by our generated queries.
-Figure 8. Achitecture of Dinkel.
-Query Reduction.
-No existing tool is available for reducing Cypher queries, which makes it difficult for developers to minimize bug-triggering queries and investigate bugs.
-To address this problem, we implemented Dinkel with an automatic query reduction method.
-The core idea of this method is to reduce queries clause by clause.
-For each Cypher clause, Dinkel first tries to delete it.
-If the query without the clause still triggers the bug, the clause will be permanently removed.
-Otherwise, the deleted clause will be recovered, and Dinkel goes on to try to replace the clause with an alternative clause if possible.
-For example, Dinkel can try to replace the CALL clause in Figure 6 with its subclause CREATE.
-If some clauses are successfully deleted or replaced in one try, Dinkel will restart the reduction process for the reduced queries.
-The process stops when no clause in the query can be further reduced.
-With this method, bug-triggering queries can be efficiently reduced.
-Bug Deduplication.
-When a query generated by Dinkel triggers a bug, we need to demonstrate whether this bug is unique or a duplicate of another bug already found by Dinkel.
-A feasible way is bisection (Git-bisect: Use binary search to find the commit that introduced a bug., 2024; Zeller, 2009), which finds the earliest bug-inducing commits for each triggered bug and deduplicates bugs by checking whether their bug-inducing commits are the same.
-As GDBMSs are large-scale system software, compiling and building a version of a GDBMS demands considerable time and resources (e.g., 8 minutes for Neo4j in our evaluation), which can make bisection infeasible when a large number of commits need to be bisected (e.g., 79K commits in Neo4j’s GitHub repository (Neo4j.
-Graphs for Everyone., 2024)).
-To improve the efficiency of bisections, Dinkel leverages containerization techniques (e.g., Docker (Docker: Accelerated Container Application Development., 2024)) to cache and manage the built GDBMS versions.
-Given a batch of bugs needed to be deduplicated, Dinkel sets up multiple containers to bisect the bug-inducing commits.
-During bisection for a bug, Dinkel needs to test many commits by checking whether the bug can be reproduced in this commit version.
-For each commit that has not been tested yet, Dinkel compiles and builds a GDBMS version, and caches the built version in a container image.
-When testing a cached commit during bisection for a new bug, Dinkel directly utilizes the image to set up the container, without repeatedly compiling and building the same versions of GDBMSs.
+Administrative clauses, such as SHOW FUNCTIONS and USE, are not supported because they are not related to query processing logic.
+For example, SHOW FUNCTIONS is designed to list all available functions and cannot be combined with other clauses.
+USE is used to switch the manipulated databases, which is not necessary as all databases are empty on initialization and they function in the same way.
+As shown in Table 4, Dinkel supports more Cypher clauses than existing approaches (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zhuang et al., 2023; Mang et al., 2024b).
+None of the existing approaches support FOREACH clauses, because this clause complicates control flow and can significantly modify graph states during query execution.
+UNION, EXISTS, and COUNT are not supported either, as these clauses invoke subqueries that can inherit and change the graph states of the main query.
+Lacking a systematic model for handling graph state changes, existing approaches are limited in generating complex queries using these clauses.
+Moreover, existing approaches support validation for only queries following specific patterns, while many clauses complicate the semantics of Cypher queries, making them unable to be validated by existing approaches.
 Table 4. Cypher clauses supported by existing approaches
@@
 COUNT
+Bug Detection.
+Dinkel detects logic bugs via its fine-grained query transformation.
+Besides logic bugs, Dinkel applies some methods to identify internal errors and memory bugs in GDBMSs.
+Specifically, for each executed query, Dinkel checks the messages returned from the GDBMS.
+If the message indicates an internal error in the GDBMS (e.g., the error in Figure 6), a bug is identified.
+We also enable all embedded assertions in the tested GDBMSs to catch assertion failures.
+In addition, we enable ASan (Address Sanitizer., 2019) for the GDBMSs developed in C/C++ to identify memory bugs triggered by our generated queries.
+Query Reduction.
+No existing tool is available for reducing Cypher queries, which makes it difficult for developers to minimize bug-triggering queries and investigate bugs.
+To address this problem, we implemented Dinkel with an automatic query reduction method.
+The core idea of this method is to reduce queries clause by clause.
+For each clause, Dinkel tries to delete it.
+If the query without the clause still triggers the bug, the clause will be permanently removed.
+Otherwise, the deleted clause will be recovered, and Dinkel goes on to try to replace the clause with an alternative clause if possible.
+For example, Dinkel can try to replace the CALL clause in Figure 2 with its subclause CREATE.
+If some clauses are successfully deleted or replaced in one try, Dinkel will restart the reduction process for the reduced queries.
+The process stops when no clause in the query can be further reduced.
 5. Evaluation
-To demonstrate the effectiveness of Dinkel in GDBMS bug detection, we evaluate it on real-world GDBMSs and seek to answer the following questions:
+Our evaluation aims to answer the following questions:
   1. Q1
-    Can Dinkel find real bugs in widely-used and extensively-tested GDBMSs? (Section 5.2)
+    Can Dinkel find real bugs in widely used and extensively tested GDBMSs? (Section 5.2)
   2. Q2
@@
   3. Q3
-    How do query context and graph schema contribute to the effectiveness of Dinkel? (Section 5.4)
+    How do different techniques contribute to the effectiveness of Dinkel? (Section 5.4)
   4. Q4
@@
 5.1. Experimental Setup
-We evaluated Dinkel on three real-world GDBMSs, Neo4j (Neo4j.
+We evaluated Dinkel on Neo4j (Neo4j.
 Graphs for Everyone., 2024), RedisGraph (RedisGraph.
-A graph database as a Redis module., 2023), and Apache AGE (Apache AGE.
-Graph database optimized for fast analysis and real-time data processing., 2024).
-We chose them because they are popular and open-source GDBMSs supporting Cypher.
-Neo4j is the most popular GDBMS, according to DB-Engines Ranking (DB-Engines Ranking., 2024), and has been extensively tested by existing approaches (Hua et al., 2023; Jiang et al., 2023b; Kamm et al., 2023; Zhuang et al., 2023).
-RedisGraph was widely adopted by Redis (the 6th most popular DBMS (DB-Engines Ranking., 2024)) to provide functionalities of graph data querying and storage.
-Its fork, FalkorDB (FalkorDB.
-A super fast Graph Database uses GraphBLAS under the hood for its sparse adjacency matrix graph representation., 2024), is also actively developed.
-Apache AGE is a widely-used extension for PostgreSQL (the 4th most popular DBMS (DB-Engines Ranking., 2024)) that enables applications to manipulate graph data on the top of a relational DBMS.
-We evaluate Dinkel on these GDBMSs with the latest versions.
-During our testing, if the code of the tested GDBMSs is updated (e.g., a new version is released), we set up new Dinkel instances to test the updated versions.
-Specifically, we test Neo4j starting from version 5.6.0, RedisGraph starting from version 2.12.0, and ApacheAGE starting from version 1.3.0.
-We clone the code of these GDBMSs from their official GitHub repositories.
-To demonstrate the effectiveness of Dinkel on testing black-box GDBMSs, we also evaluate Dinkel on the Enterprise version of Neo4j from 5.6.0.
+A graph database as a Redis module., 2023), and Memgraph (Memgraph.
+Open-source graph database, tuned for dynamic analytics environments.
+Easy to adopt, scale and own., 2024), which are popular and extensively tested by existing approaches (Hua et al., 2023; Jiang et al., 2023b; Kamm et al., 2023; Zhuang et al., 2023; Mang et al., 2024b).
+We evaluate Dinkel on the latest GDBMS versions.
+During testing, if the tested GDBMSs are updated (e.g., a new version is released), we set up new Dinkel instances to test the updated versions.
+Specifically, we test Neo4j from version 5.6.0, RedisGraph from 2.12.0, and Memgraph from 2.7.0.
+RedisGraph is no longer maintained after 2.12.10, and thus we move to test its successor, FalkorDB (FalkorDB.
+A super fast Graph Database uses GraphBLAS under the hood for its sparse adjacency matrix graph representation., 2024), from its first release version 4.0.0.
+We count the bugs of RedisGraph and FalkorDB together as they share the majority of their code base.
+To show its effectiveness for black-box testing GDBMSs, we evaluate Dinkel on the Enterprise version of Neo4j from 5.6.0 onward.
 Each time we implement new features on Dinkel, we stop and restart the testing.
-Overall, the testing process lasts three months.
-The evaluation was performed on Ubuntu 20.04 with a 64-core AMD EPYC 7742 processor running at 2.25GHz and 256GB of RAM.
+Overall, the testing campaign intermittently lasted 9 months.
+The evaluation was performed on a machine running Ubuntu 20.04, with a 64-core AMD EPYC 7742 processor running at 2.25GHz and 256GB of RAM.
 5.2. Bug Detection
+Table 5 shows the status of the bugs found by Dinkel.
+In total, Dinkel found 127 unique bugs, including 51 bugs in Neo4j, 62 in RedisGraph, and 14 in Memgraph.
+Among these bugs, 113 are confirmed, and 84 are fixed.
+None of the 127 bugs are duplicates.
+As mature GDBMSs, Neo4j, RedisGraph, and Memgraph have been extensively tested in both industry and academia.
+The significant number of new bugs demonstrates the powerful bug-finding ability of Dinkel.
+We will publish all our 127 bug reports afterward with our artifact.
+Bug Classification. We classify the bugs found by Dinkel into three categories according to their manifestation:
+  • Logic bugs.
+The tested GDBMSs produce incorrect results for specific queries.
+These bugs were exposed because they lead to discrepancies between the results of the original queries and the transformed queries generated by Dinkel.
+  • Internal errors.
+The tested GDBMSs unexpectedly throw exceptions or errors when processing syntactically and semantically valid queries.
+The error messages can indicate the inconsistency of the internal execution status.
+  • Crashes. The queries cause GDBMSs to crash due to assertion failures or memory corruption.
 Table 5. Status of the bugs found by Dinkel
 GDBMS Reported Confirmed Fixed
-Neo4j 32 32 32
-RedisGraph 17 17 10
-Apache AGE 11 9 9
-Total 60 58 51
-Table 5 shows the status of the bugs found by Dinkel.
-In total, Dinkel found 60 unique bugs, including 32 bugs in Neo4j, 17 in RedisGraph, and 11 in Apache AGE.
-Among these bugs, 58 are confirmed, and 51 are fixed.
-None of the 60 bugs are duplicates.
-Bug Classification. We classify the bugs found by Dinkel into two categories according to their manifestation:
-  • Internal errors.
-The tested GDBMSs unexpectedly throw exceptions or errors when processing syntactically and semantically valid queries.
-The error messages can indicate the inconsistency of the internal execution status (e.g., the error triggered by the query shown in Figure 6).
-  • Crashes. The queries cause GDBMSs to crash due to assertion failures or memory corruption, e.g., a null-pointer dereference.
+Neo4j 51 51 49
+RedisGraph 62 48 29
+Memgraph 14 14 6
+Total 127 113 84
 Table 6. Classifying the found bugs
-GDBMS Internal Error Crash
-Neo4j 32 0
-RedisGraph 3 14
-Apache AGE 8 3
-Total 43 17
-Note that Neo4j is realized in Java and implements exception handling for unexpected errors (e.g., null dereferences in Figure 6 and out-of-bound accesses in Figure 16), and thus Neo4j does not crash but posts exception information when triggering memory corruption.
-Therefore, we classify all the bugs Dinkel found in Neo4j as internal errors.
+GDBMS Logic Bug Internal Error Crash
+Neo4j 13 38 0
+RedisGraph 18 10 34
+Memgraph 2 8 4
+Total 33 56 38
 Table 6 shows the classification results.
-Among the 60 Dinkel found, 43 cause internal errors, 32 of which are Neo4j bugs.
-Dinkel can find 17 bugs that crash RedisGraph and Apache AGE, which are implemented in C/C++.
-Among these 17 bugs, 8 are caused by memory corruptions, and 9 are caused by assertion failures.
-These results demonstrate that Dinkel is effective in finding bugs in GDBMSs.
+Among the 127 bugs Dinkel found, 33 are logic bugs, 56 cause internal errors, and 38 result in GDBMS crashes.
+Note that Neo4j is implemented in Java and has exception handling for unexpected memory errors.
+Thus, Neo4j does not crash, but rather responds with exception information when triggering memory corruption.
+Therefore, Dinkel found no crashes in Neo4j, but 38 internal errors.
+Among the 38 crashes found in RedisGraph and Memgraph, 29 are caused by memory corruptions, and 9 are caused by assertion failures.
+These results demonstrate that Dinkel can comprehensively test GDBMSs by finding various bugs.
+Figure 6 shows a bug that triggers a Neo4j internal error.
+This query constructs an empty array [] and invokes a subquery using a CALL.
+The subquery also constructs an empty array.
+It then invokes two UNWIND clauses, which iterate over each element in the operated array.
+For each iterated element, the UNWIND executes the subsequent clause under the context of this element.
+For example, the array [x] used by the second UNWIND references the variable x, which is 0 when the first UNWIND iterates over the item 0 in the array [0].
+After the CALL, the query utilizes FOREACH, whose execution depends on the query context of the CALL.
+To optimize query execution, Neo4j tries to flatten the FOREACH loop.
+However, such optimization does not work well when the loop involves update operations (i.e., MERGE) under complicated contexts (i.e., the contexts produced by CALL).
+The improper optimization corrupts the internal data structures of Neo4j, Eagers, which are the production of another optimization, Eagerness analysis.
+In the end, an internal error is triggered when Neo4j tries to access the corrupted Eagers.
+To fix this bug, Neo4j developers modify both the Eagerness analysis and the flattening strategy for FOREACH clauses to ensure they work consistently.
 Affected GDBMS Components.
-We investigated 42 of the 51 fixed bugs, where we can analyze the fix patches to identify the GDBMS components affected by the bugs accurately.
-The 9 fixed bugs in the Enterprise version of Neo4j are not included because developers keep the fix patches confidential.
+We investigated 59 of the 84 fixed bugs, where we were able to analyze the fix patches to identify the GDBMS components affected by the bugs accurately.
+The 19 fix patches in the Enterprise version of Neo4j are confidential and thus not included.
+The 6 fixed bugs in Memgraph are also not included because Memgraph developers mix the fixes into their regular development commits, where we cannot accurately analyze the affected components.
 Table 7 shows the results.
-Among the 42 fixed bugs, 32 affect the parsers of GDBMSs, 15 affect the planners, and 10 affect the executor.
-Interestingly, 9 bugs can affect two components, and 4 can affect all the components.
-For example, the bug shown in Figure 16 can affect both the parser and planner of Neo4j.
-In total, 19 bugs can affect either the planner, the executor, or both.
-These results demonstrate that (1) Dinkel can find bugs in various GDBMS components; and (2) Dinkel can find bugs in the deep logic of GDBMS implementation, considering 45% (19/42) of bugs are related to the planners or executors.
-Table 7. GDBMS components affected by bugs
+Among the 59 fixed bugs, 29 affect the parsers of GDBMSs, 25 affect the planners, and 20 affect the executor.
+Interestingly, 29 bugs (5 logic bugs) affect two components, and 12 bugs (1 logic bug) affect all the components.
+For example, the bug shown in Figure 10 affects both the planner and executor of Neo4j.
+In total, 40 bugs (8 logic bugs) affect either the planner, the executor, or both.
+The 6 logic bugs affecting the executor also affect the planner.
+These results demonstrate that (1) Dinkel can find bugs in various GDBMS components; and (2) Dinkel can find bugs in the deep logic of GDBMSs, considering 67.8% (40/59) of fixed bugs are related to the planners or executors.
+WITH [] AS n0 ORDER BY null
+CALL {
+WITH [] AS n1 ORDER BY null
+UNWIND [0] AS x
+UNWIND [x] AS n2
+RETURN n2 AS n3
+} FOREACH (n4 IN null | MERGE ({key:n3}))
+Figure 6. A query triggering a Neo4j internal error—Neo4j- Error: ExecutionFailed (arraycopy: last destination index 6 out of bounds for object array[5]).
+Table 7. GDBMS components affected by bugs. The number in the parentheses is for the logic bugs
 GDBMS Parser Planner Executor
-Neo4j 17 11 6
-RedisGraph 7 4 4
-Apache AGE 8 0 1
-Total 32 15 11
-Diversity of Cypher Features.
-We investigate the queries triggering the 60 bugs and analyze the distribution of the Cypher features used in these queries.
-All these queries are reduced.
+Neo4j 18 (0) 18 (5) 10 (3)
+RedisGraph 15 (1) 15 (3) 19 (3)
+Total 23 (1) 23 (8) 29 (6)
+Data Dependencies for Triggering Bugs.
+To convey the data dependencies required to trigger bugs, we analyze dependencies in the 127 bug-triggering queries.
 The results are shown in Figure 7.
-The RETURN keyword is present in most bug-triggering queries, appearing in 64% of them.
-This is unsurprising, as RETURN is required to appear at the end of a query containing only reading clauses (e.g., MATCH), otherwise, the query would be syntactically incorrect.
-MERGE (33%) and CREATE (26%) are the two clauses inducing the creation of graph properties (e.g., providing available elements for the operations of subsequent clauses).
-We can deduce from this that graph data is often required to be present for bugs to trigger.
-The remaining keywords can induce complicated data flow (e.g., WITH (22%) and UNWIND (14%)) or control flow (e.g., CALL (14%) and CASE WHEN (14%)) in the query, indicating that bugs often appear when the database system has to handle such flows.
-ORDER BY also appears frequently (14%), which forces GDBMSs to keep track of the manipulated graph data and perform non-trivial comparisons to ensure correct order.
-Figure 9. Feature distribution of bug-triggering queries.
-Data Dependencies for Triggering Bugs.
-To convey the query complexity Dinkel required to trigger bugs, we analyze the complexity of the 60 bug-triggering queries.
-The results are shown in Figure 10.
-Nearly half of the bugs (45%) require at least one data dependency.
-Specifically, 13 bug-triggering queries contain one data dependency, 8 contain two, and 6 contain three or more.
-For example, in Figure 3, the bug-triggering query contains 8 data dependencies, as illustrated in Figure 11.
+Over half of the bugs (58%) require at least one data dependency.
+Specifically, 42 bug-triggering queries contain one data dependency, 16 contain two, 11 contain three, and 7 contain four or more.
+For example, in Figure 1, the query contains 8 data dependencies, as illustrated in Figure 8.
 Among the 8 dependencies, 6 are involved in query context (i.e., the variables x and y assigned to a node and a relationship), and 2 are involved in graph schema (i.e., the label A).
@@
 Figure 8. Data dependencies within the query in Figure 1.
-Figure 12. Query sizes of bug-triggering queries.
 Size of Bug-Triggering Queries.
-Figure 12 shows the size of the 60 bug-triggering queries.
-54 bugs can be triggered by queries whose size is less than 100 bytes.
-The bug-triggering queries shown in Figure 3 and Figure 6 are examples of such queries.
-As query size increases from 0 to 100 bytes, the number of triggered bugs increases almost linearly.
-The bug-triggering query with the largest size is 146 bytes and is shown in Figure 16.
-This query triggers an internal error caused by an out-of-bounds access in Neo4j.
+Figure 9 shows the size of the 127 bug-triggering queries.
+All these queries have been reduced using the methods mentioned in Section 4.
+112 bugs can be triggered by queries whose size is less than 120 bytes.
+The bug-triggering query shown in Figure 1 is an example of such a query.
+As query size increases from 0 to 120 bytes, the number of triggered bugs increases almost linearly.
+The bug-triggering query with the largest size is 249 bytes, which triggers an internal error in Neo4j and was fixed by their developers.
 Bug Importance.
@@
 The bugs in the Enterprise version are critical because this version is commonly deployed on commercial applications.
-Among the 32 Neo4j bugs Dinkel found, 19 bugs can be triggered in both Enterprise and Community versions, and 8 bugs can be triggered in only the Enterprise version.
-RedisGraph and Apache AGE did not provide severity information about the reported bugs, but we noticed that most of the bugs were fixed, which indicates that the bugs are non-trivial.
-Some developers express their appreciation for our effort in finding bugs in their GDBMSs. Particularly, Neo4j provides very positive feedback written in German. We translate it into English:
+Among the 51 Neo4j bugs Dinkel found, 32 bugs can be triggered in both Enterprise and Community versions, and 19 bugs can be triggered in only the Enterprise version.
+RedisGraph and Memgraph did not provide severity information about the reported bugs, but we noticed that most bugs were fixed, indicating that the bugs are non-trivial.
+Some developers express their appreciation for our effort in finding bugs in their GDBMSs.
+Particularly, Neo4j provides very positive feedback:
 My colleague has told me that you’ve been busy creating GitHub issues and have contributed to improving Neo4j! Thanks a lot for that! In the name of Neo4j, I would like to send you some swag […]
 5.3. Query Generation
-To understand the complexity and validity of the queries generated by Dinkel, we ran Dinkel on each tested GDBMS for 24 hours and collected all generated queries.
+To understand the complexity and validity of the queries constructed by Dinkel, we ran Dinkel on each tested GDBMS for 48 hours and collected all generated queries (transformed queries are not included).
 Table 8 shows the statistical results.
-Table 8. Queries generated by Dinkel within 24 hours
-GDBMS Gen. Valid Clause Dep. Size
-Neo4j 68K 89.88% 23.68 30.16 1202.39
-RedisGraph 956K 98.14% 23.42 30.15 992.41
-Apache AGE 367K 81.79% 16.94 15.60 830.35
-Total 1391k 93.43% 21.72 26.31 959.85
-  • Gen.: the number of generated queries;
-    Valid: the number of valid queries among the generated ones;
-    Clause, Dep., and Size: the average number of clauses, data dependencies, and bytes over all generated queries, respectively.
 Query Complexity.
-For the 1391K generated queries, Dinkel constructs 30.23M clauses.
-The average number of clauses for each query is 21.72.
-Additionally, on average, each query contains 26.31 data dependencies.
-Because of the large number of query clauses and data dependencies, the queries generated by Dinkel are typically big.
-The average query size is 959.85 bytes.
+On average, Dinkel generates 655k queries for each GDBMS, with 22.62M clauses constructed.
+The average number of clauses of each query is 34.54, and the number of data dependencies is 26.36.
+Because of the large number of clauses and dependencies, the queries generated by Dinkel are typically large.
+The average query size is 1153.3 bytes.
 These results demonstrate that Dinkel can effectively generate complex queries, which are large and contain multiple clauses with complicated data dependencies.
 The generated queries are more complex than the bug-triggering queries discussed in Section 5.2, because all the bug-triggering queries are reduced, while the generated queries inevitably contain many redundant parts (Jiang et al., 2023c; Maciver and Donaldson, 2020; Regehr et al., 2012).
-We find that the queries generated for Apache AGE are less complex (i.e., fewer clauses, fewer data dependencies, and smaller size), because Apache AGE supports fewer clauses, and many clauses that can induce data dependencies are missing.
-As a result, the relative chance of Dinkel choosing clauses (e.g., RETURN) that terminate query generation becomes much higher, and thus the queries generated for Apache AGE exhibit lower complexity.
+Figure 9. Query sizes of bug-triggering queries.
+Table 8.
+Queries generated by Dinkel within 48 hours.
+Gen.: the number of generated queries; Valid: the number of valid queries; Clause, Dep., and Size: the average of clauses, data dependencies, and bytes over all queries, respectively
+GDBMS Gen. Valid Clause Dep. Size
+Neo4j 318k 89.90% 40.97 26.77 1394.69
+RedisGraph 1’078k 93.90% 32.71 30.59 1071.58
+Memgraph 570k 83.27% 29.95 21.71 993.63
+Average 655k 89.02% 34.54 26.36 1153.3
 Query Validity.
-Among the 1391K Cypher queries generated by Dinkel, 1300K are valid.
-The percentage of valid queries is 93.43%.
+Among the 655k queries generated by Dinkel, 583k are valid.
+The percentage of valid queries is 89.02%.
 The result demonstrates that Dinkel can keep a high validity rate even when generating complex queries.
 We investigated the invalid queries generated by Dinkel and found that they were mainly caused by illegal arithmetic operations.
-Some arithmetic operations require their operands to satisfy some constraints (e.g., the divisor in a division operation must be non-zero).
+Some arithmetic operations require the operands to satisfy some constraints (e.g., the divisor in a division must be non-zero).
 However, Dinkel cannot track the value of each operator in queries because the value may depend on complex expressions, whose results are difficult to calculate (e.g., hash functions).
-Queries failing to satisfy the constraints of operations cause semantic errors like divisions by zero.
-We noticed that the validity of queries generated for Apache AGE is lower than that of the other two GDBMSs.
-This is caused by the fact that Apache AGE is designed to be a PostgreSQL extension and thus needs to interact with the SQL contexts of PostgreSQL.
-Specifically, the columns returned by Cypher queries of Apache AGE have to match the type definitions in the wrapping SQL queries used by PostgreSQL, otherwise, the SQL queries become invalid.
-Figure 13 shows an example, where the Cypher query is valid but the overall SQL query is invalid because the Cypher query returns two columns while the SQL query expects only one column.
-Such invalid cases do not affect bug detection, because the error of SQL queries occurs only after the Cypher queries are fully executed.
-For example, in Figure 13, even though the SQL query is invalid, it can crash Apache AGE and PostgreSQL, because the internal Cypher query triggers the bug before it returns values to the SQL query.
-1SELECT * FROM cypher('graph',\$\$
-2 CREATE (x) SET x.n0 = (true OR true)
-3 RETURN 0 AS y, 1 AS z
-4\$\$) AS (v agtype);
-Figure 13. An invalid SQL query that wraps a valid Cypher query, which crashes Apache AGE and PostgreSQL.
+Queries failing to satisfy the constraints of operations cause semantic errors, such as divisions by zero.
 Throughput.
-On average, one Dinkel instance generates 5.36 test cases per second.
-Specifically, at each second, one Dinkel instance generates 0.78 test cases for Neo4j, 11 test cases for RedisGraph, and 4.25 test cases for Apache AGE.
-To understand the bottleneck of test throughput, we further investigated the time used by Neo4j in query generation and query execution, and found that 92.79% of CPU time is occupied by Neo4j executing the generated queries, indicating that the testing throughput is determined by the performance of the tested GDBMS.
+On average, one Dinkel instance generates 3.79 tests (each test consists of one original query and one transformed query) per second.
+Specifically, at each second, one Dinkel instance completes 1.84 tests for Neo4j, 6.24 for RedisGraph, and 3.30 for Memgraph.
+To understand the bottleneck of test throughput, we further investigated the time used by Neo4j in query generation, transformation, and execution.
+We found that for Neo4j, 74.75% of CPU time was used to execute the generated queries, which are complex and can be time-consuming.
+This result indicates that the test throughput is dominated by the performance of the tested GDBMSs.
 We believe the current throughput is practical considering (1) GDBMS testing typically lasts for several months (Kamm et al., 2023; Jiang et al., 2023b), and thus a sufficient number of test cases can be executed; and (2) setting up multiple Dinkel instances can significantly improve the test efficiency.
-5.4. Ablation Analysis
-To understand the contribution of query context and graph schema for bug detection in GDBMSs, we analyze the 60 bug-triggering queries by investigating whether the generation for these queries depends on specific graph states.
-Specifically, we extract the data dependencies contained by each query and check whether at least one of the data dependencies is related to query context or graph schema.
-Table 9 shows the analysis results.
-Among the 60 bug-triggering queries, 22 depend on only query context, 1 depends on only graph schema, and 4 depend on both (e.g., the bug-triggering query shown in Figure 11).
-According to these results, without query context, 26 bugs cannot be found.
-This demonstrates that query context is important because it enables Dinkel to generate queries referencing specific nodes, relationships, or expressions constructed in internal clauses.
-Without graph schema, a further 5 bugs cannot be found.
+5.4. Importance of Key Techniques
+To show the importance of our techniques, we analyze how they contribute to the 127 bugs found by Dinkel.
+The analysis consists of two parts.
+The first focuses on the query context and graph schema in query generation, while the second assesses the transformation strategies (4 clause-level ones and 2 expression-level ones) used in test-oracle construction.
+We analyzed the 127 bug-triggering queries and checked whether their generations depend on specific graph states.
+Specifically, we extracted the data dependencies contained by each query and check whether at least one dependency is related to query context or graph schema.
+Table 9 shows the results.
+Among the 127 queries, 72 depend on only query context, 5 depends on only graph schema, and 9 depend on both.
+According to these results, without query context, 81 bugs (12 logic bugs) cannot be found.
+This demonstrates the importance of query contexts, which enable Dinkel to generate queries referencing specific nodes, relationships, or expressions constructed in internal clauses.
+Without graph schema, 14 bugs (3 logic bugs) will be missed.
 These bug-triggering queries require Dinkel to properly reference the latest graph labels or properties.
-The following shows two queries involving query context and graph schema, respectively.
-Table 9. Bug-triggering queries related to graph states
+Table 9. Bug-triggering queries related to QC: only query context; GS: only graph schema; Both: both graph states. The number in the parentheses is for the logic bugs
 Version Found QC GS Both
-Neo4j 32 11 1 2
-RedisGraph 17 6 0 2
-Apache AGE 11 5 0 0
-Total 60 22 1 4
-  • QC: the queries depend on only query context;
-    GS: the queries depend on only graph schema;
-    Both: the queries depend on two graph states.
-Example 1: Query referencing query context.
-The query shown in Figure 14 triggers an internal error.
-This query uses a UNWIND clause to iterate over the elements in the array [0]+[] (i.e., [0]).
-Then, the query invokes toBoolean to convert the iterated element to the boolean type and return each converted element.
-However, when concatenating the arrays [0] and [], Neo4j incorrectly maintains the type information of the concatenated array of [0] and [].
-It directly assigns the type of elements in the right-hand-side array (i.e., []), which is implicit and not determined, to elements of the concatenated array.
-In the end, the type of the elements in the concatenated array is unexpectedly assigned as Float (while Integer is the expected type), resulting in the internal error of type mismatching.
-To fix this bug, when concatenating two arrays, developers make Neo4j consider the element type in both arrays.
-Example 2: Query referencing graph schema.
-In Figure 15, the query specifies a graph pattern ()-[:A]->({x:0}), where the relationships with type A are connecting an arbitrary node to another node with property x equal to 0.
-The MERGE clause tries to find the nodes and relationships satisfying the pattern.
-If there is no such node and relationship, the MERGE clause creates new ones satisfying the patterns.
-After the MERGE clause, the EXISTS clause checks whether the graph pattern (:!(B&C))-[{x:0}]->() can be matched in the current graph database and returns the corresponding boolean value.
-The EXISTS clause references two non-existing labels, B and C, and the property x created in the MERGE clause.
-As the bug is triggered in the Enterprise version of Neo4j, the root cause and the corresponding fixing patch are kept confidential by Neo4j developers.
-1UNWIND [0]+[] AS i
-2RETURN toBoolean(i)
-Figure 14. A query that references query context and triggers a Neo4j internal error—Type mismatch: expected Boolean, Integer or String but was Float (line 1, column 37 (offset: 36))
-1MERGE ()-[:A]->({x:0})
-2RETURN EXISTS {(:!(B&C))-[{x:0}]->()};
-Figure 15.
-A query that references graph schema and triggers a Neo4j internal error—Neo4jError: ExecutionFailed (Cannot invoke ”RelationshipDataAccessor.properties(…)” because ”relationshipCursor” is null).
+Neo4j 51 (13) 23 (3) 3 (0) 5 (2)
+RedisGraph 62 (18) 34 (7) 1 (1) 4 (0)
+Memgraph 14 (2) 5 (0) 1 (0) 0 (0)
+Total 127 (33) 72 (10) 5 (1) 9 (2)
+Table 10.
+Transformation strategies used in finding logic bugs, including Dead Code (DC), Unused Context (UC), Redundant Write (RW), Inconsequential Supersession (IS), Mathematical Identity (MI), and Cypher Feature (CF)
+Type Strategy Neo4j RedisGraph Memgraph Total
+Clause DC 3 4 0 7
+UC 2 3 0 5
+RW 1 3 1 5
+IS 2 6 0 8
+Total 8 16 1 25
+Expr MI 1 1 1 3
+CF 4 1 0 5
+Total 5 2 1 8
+Total 13 18 2 33
+We investigated the 33 logic bugs found by Dinkel, and checked whether their bug-triggering queries use specific transformation strategies.
+Table 10 shows the results.
+We find that each logic bug was triggered by using just one transformation rule in Table 3.2 and Table 3.2.1.
+Specifically, among the 33 logic bugs, 25 are exposed by clause-level transformations, and 8 by expression-level transformations.
+The number of bugs found by each transformation strategy is close to the average (i.e., 5.5).
+Inconsequential Supersession finds the most bugs with 8, and Mathematical Identity finds the least with 3 bugs.
+These results indicate that all strategies are effective in identifying logic bugs in GDBMSs.
+Moreover, the 25 bugs found by clause-level transformations demonstrate the effectiveness of our fine-grained transformation over EET (Jiang and Su, 2024), which supports only expression-level transformation.
+The following discusses two representative bug examples.
+Example 1: Figure 10 shows a Neo4j bug that causes an incorrect query result.
+The original query references both query contexts and graph schema.
+Specifically, the first CREATE clause creates two nodes with a relationship y, labeled A.
+One node has a property x with value 0.
+The second CREATE clause creates a new node with a property x whose value is equal to the property x of the end node of y (i.e., 0).
+The RETURN clause returns the number of nodes whose property x is 0.
+The expected result should be 2.
+However, Neo4j produces 1, because of a logic bug triggered when Neo4j processes two consecutive CREATEs that have data dependencies.
+Dinkel finds this bug using a clause-level transformation, Redundant Write.
+It inserts a CREATE clause and DELETE clause between the two CREATEs.
+Such a transformation preserves the query semantics but makes the two CREATEs not consecutive anymore, and thus Neo4j produces a correct result for the transformed query.
+The developers fixed this bug by significantly reconstructing the source code related to CREATE clauses.
+In the end, 17 source files were modified.
+// Original: 1
+CREATE ({x:0})<-[y:A]-()
+CREATE ({x:endNode(y).x})
+RETURN COUNT {({x:0})} AS n
+// Transformed: 2
+CREATE ({x:0})<-[y:A]-()
+CREATE (z) DELETE z
+CREATE ({x:endNode(y).x})
+RETURN COUNT {({x:0})} AS n
+Figure 10. Clause-level transformation reveals a bug.
+Example 2: Figure 11 shows a logic bug in Neo4j found by expression-level transformation, Mathematical Identity.
+It is based on a rule that the result of NULL XOR a is NULL, no matter what a is.
+Dinkel transforms the NULL into NULL XOR EXISTS{()}, expecting that the result should remain NULL.
+However, Neo4j produces FALSE.
+The root cause of this bug is related to an operator in the Cypher planner of Neo4j, LetSelectOrSemiApply.
+It is invoked only when the transformed query is processed.
+Such an operator fails to work correctly because it does not consider NULL expressions in its arguments.
+To fix this bug, Neo4j developers refined this operator by adding additional logic to check and handle NULL in its arguments.
+// Original: NULL
+RETURN NULL AS x
+// Transformed: FALSE
+RETURN NULL XOR EXISTS{()} AS x
+Figure 11. Expression-level transformation reveals a bug.
 5.5. Comparison
 Bug Latency Study.
-To demonstrate that Dinkel can find bugs missed by existing approaches (Kamm et al., 2023; Hua et al., 2023; Jiang et al., 2023b; Zhuang et al., 2023; Mang et al., 2024), we investigate the latencies of the bugs found by Dinkel.
+To demonstrate that Dinkel can find bugs missed by existing approaches, we investigate the latencies of bugs found by Dinkel.
 For each bug, we check whether its bug-inducing commit was created before the years when existing approaches were published.
 If Dinkel finds some long-latent bugs, we can conclude that Dinkel can find bugs missed by existing approaches.
-This comparison is reasonable and objective because: (1) none of the bugs found by Dinkel are marked as duplicated by GDBMS developers, which means that no approach found these bugs until Dinkel found them; and (2) all existing approaches have extensively tested Neo4j and RedisGraph (Kamm et al., 2023; Hua et al., 2023; Jiang et al., 2023b; Zhuang et al., 2023; Mang et al., 2024), which means that in these two GDBMSs, no approach found the long-latent bugs found by Dinkel during their evaluation.
-GDsmith (Hua et al., 2023), GDBMeter (Kamm et al., 2023), GraphGenie (Jiang et al., 2023b), and GAMERA (Zhuang et al., 2023) were published in 2023, and GRev (Mang et al., 2024) was published in 2024.
+This comparison is reasonable and objective because: (1) none of the bugs found by Dinkel are marked as duplicated by developers, meaning that no approach found these bugs until Dinkel found them; and (2) all existing approaches have extensively tested Neo4j and RedisGraph (Kamm et al., 2023; Hua et al., 2023; Jiang et al., 2023b; Zhuang et al., 2023; Mang et al., 2024b), meaning that in these two GDBMSs, no approach found the long-latent bugs found by Dinkel during their evaluation.
+Some existing approaches do not test Memgraph, but we still include it to show the effectiveness of Dinkel.
+GDsmith (Hua et al., 2023), GDBMeter (Kamm et al., 2023), GraphGenie (Jiang et al., 2023b), and GAMERA (Zhuang et al., 2023) were published in 2023, and GRev (Mang et al., 2024b) was published in 2024.
 Therefore, for each bug, we checked whether its bug-inducing commit was created before 2024 (i.e., in 2023 or earlier) and 2023.
-Table 10 shows the results.
-Among the 49 bugs Dinkel found in Neo4j and RedisGraph, 33 bugs already existed before 2023, indicating that all the existing approaches proposed after 2023 failed to find these 33 bugs in their extensive evaluation.
-In addition, 14 bugs were introduced to Neo4j and RedisGraph between 2023 and 2024, while GRev, which was proposed after 2024, cannot find these bugs.
+As shown in Table 11, among the 113 bugs in Neo4j and RedisGraph, 47 already existed before 2023, among which 10 are logic bugs.
+It indicates that all existing approaches failed to find these 47 bugs in their extensive evaluation.
+In addition, 45 bugs (24 logic bugs) were introduced to Neo4j and RedisGraph between 2023 and 2024, while GRev, which was proposed after 2024, cannot find these bugs.
+In addition, GDsmith and GRev support Memgraph, but neither can find the 4 long-latent bugs induced before 2023.
 These results indicate that existing approaches indeed miss many long-latent bugs, while Dinkel can effectively find them.
-Table 10. Latency of the logic bugs found by Dinkel
+Table 11. Latency of the bugs found by Dinkel. The number in the parentheses is for the logic bugs
 DBMS Found Bug-involved year Longest latency
 ¡ 2024 ¡ 2023
-Neo4j 32 30 20 2016
-RedisGraph 17 17 13 2020
-Apache AGE 11 11 7 2021
-Total 60 58 40 2016
-We analyze the bug-triggering queries of these bugs and conclude the reason why existing approaches missed these bugs as the following: (1) the bug-triggering queries (e.g., the queries shown in Figure 3 and Figure 6) contain data dependencies, which existing approaches lack support for handling; and (2) the bug-triggering queries contain Cypher features (e.g., CALL and FOREACH in the query shown in Figure 16) that are not supported by existing approaches.
-The following is an example of the 33 bugs.
-1WITH [] AS n0 ORDER BY null
-2CALL {
-3 WITH [] AS n1 ORDER BY null
-4 UNWIND [0] AS x
-5 UNWIND [x] AS n2
-6 RETURN n2 AS n3
-7}
-8FOREACH (n4 IN null | MERGE ({key:n3}))
-Figure 16. A query triggering a Neo4j internal error—Neo4j-
-Error: ExecutionFailed (arraycopy: last destination index 6 out of bounds for object array[5]).
-Example: Query using data dependencies and advanced Cypher features.
-Figure 16 shows a bug in Neo4j, which is present in 2022 but none of the existing approaches found it.
-The bug-triggering query involves complicated data dependencies and advanced Cypher clauses (e.g., FOREACH and UNWIND).
-This query constructs an empty array [] and invokes a subquery using the CALL clause.
-The subquery also constructs an empty array.
-It then invokes two UNWIND clauses, which iterate over each element in the operated array.
-For each iterated element, the UNWIND executes the subsequent clause under the context of this element.
-For example, the array [x] used by the second UNWIND references the variable x, which is 0 when the first UNWIND iterates over the item 0 in the array [0].
-After the CALL, the query utilizes FOREACH, whose execution depends on the query context of the subquery in the CALL.
-To optimize the query execution, Neo4j tries to flatten the FOREACH loop.
-However, such optimization does not work well when the loop involves update operations (i.e., MERGE) under complicated contexts (i.e., the contexts produced by CALL).
-The improper optimization corrupts the internal data structures of Neo4j, Eagers, which are the production of another optimization, Eagerness analysis.
-In the end, an internal error is triggered when Neo4j tries to access the corrupted Eagers.
-To fix this bug, Neo4j developers modify both the Eagerness analysis and the flattening strategy for FOREACH clauses to ensure they work consistently.
+Neo4j 51 (13) 40 (9) 27 (7) 2016
+RedisGraph 62 (18) 52 (15) 20 (3) 2020
+Memgraph 14 (2) 9 (2) 4 (0) 2021
+Total 127 (33) 101 (26) 51 (10) 2016
+Table 12. The clauses used by bug-triggering queries. The number in the parentheses is for the logic bugs. N/A indicates that the GDBMS does not support the clause
+Clause Neo4j RedisGraph Memgraph Total
+MATCH 20 (7) 28 (6) 5 (1) 53 (14)
+CREATE 17 (9) 26 (11) 2 (1) 55 (21)
+MERGE 12 (4) 31 (12) 5 (1) 48 (17)
+DELETE 4 (4) 16 (6) 1 (1) 21 (11)
+REMOVE 0 1 (0) 0 1 (0)
+SET 4 (2) 3 (1) 0 7 (3)
+UNWIND 9 (1) 11 (6) 1 (0) 21 (7)
+WITH 16 (3) 23 (8) 0 39 (11)
+RETURN 40 (8) 39 (13) 11 (1) 90 (22)
+CALL 15 (3) 22 (8) 4 (0) 41 (11)
+FOREACH 8 (5) 6 (3) 0 14 (8)
+UNION 10 (3) 7 (3) 1 (0) 18 (6)
+EXISTS 10 (3) N/A N/A 10 (3)
+COUNT 10 (4) N/A N/A 10 (4)
+Clause Analysis.
+We analyze the clauses used in the 127 bug-triggering queries and show the results in Table 12.
+The results indicate that the bugs found by Dinkel are related to various Cypher clauses.
+Combining Table 4 and Table 12, we can demonstrate that many bugs found by Dinkel cannot be found by existing approaches, because they do not even support the clauses used in some bug-triggering queries (e.g., 14 bugs (8 logic bugs) related to FOREACH clauses).
+As discussed in Section 4, existing approaches cannot support these clauses because they lack a systematic model for query generation and a general test oracle for query validation.
 Empirical Comparison.
-To empirically demonstrate the effectiveness of Dinkel on code coverage and bug detection over the state-of-the-art, we evaluated Dinkel and existing Cypher testing tools (i.e., GDSmith, GDBMeter, GraphGenie, GAMERA, and GRev) on Neo4j and RedisGraph, which are the only two GDBMSs supported by all these tools.
-Each evaluation persisted for 48 hours.
-Table 11 shows the comparison results of Dinkel and existing tools.
-Table 11. Results of comparison
+To empirically demonstrate that Dinkel outperforms the state-of-the-art on code coverage and bug detection, we further evaluated Dinkel and existing approaches on Neo4j and RedisGraph, which are the only two GDBMSs supported by all these approaches.
+Each evaluation persisted for 48 hours and was repeated 5 times.
+Table 13 shows the comparison results.
+Table 13. Results of covered lines (average $\pm$ standard deviation) and found bugs of existing approaches over 5 runs
 Neo4j RedisGraph
-Covered Line Bug Covered Line Bug
-GDsmith 533k 0 15k 0
-GDBMeter 283k 1 19k 1
-GraphGenie 506k 0 3k 0
-GAMERA 283k 1 3k 0
-GRev 601k 1 14k 1
-Dinkel 737k 11 20k 18
-Figure 17. Covered lines of Neo4j and RedisGrpah.
+Line (Avg $\pm$ SD) Bug Line (Avg $\pm$ SD) Bug
+Dinkel 750k $\pm$ 5.2k 19 20k $\pm$ 0.1k 26
+GDBMeter 283k $\pm$ 0.2k 1 19k $\pm$ 0.3k 1
+GRev 597k $\pm$ 4.0k 1 14k $\pm$ 0.1k 1
+GAMERA 283k $\pm$ 0.1k 1 3k $\pm$ 0.1k 0
+GDsmith 515k $\pm$ 8.6k 0 15k $\pm$ 0.1k 0
+GraphGenie 506k $\pm$ 2.0k 0 3k $\pm$ 0.1k 0
 Code Coverage.
-As shown in Table 11, on average, Dinkel can cover 67% and 85% more code than existing approaches in Neo4j and RedisGraph, respectively.
-Among the tools except for Dinkel, GRev covers the most code (i.e., 601k lines) in Neo4j.
-Compared to GRev, Dinkel covers 23% more code.
-Benefiting from the powerful capability of generating complex Cypher queries, Dinkel can cover much deeper logic of GDBMSs that is related to processing advanced Cypher features and complicated data dependencies.
-In contrast, simple queries generated by existing tools have little chance to touch such logic.
-In RedisGraph, except for Dinkel, GDBMeter covers the most code (i.e., 19k lines).
-Compared to GDBMeter, Dinkel covers 5% more code.
+As shown in Table 13, on average, Dinkel covers 71% and 85% more code than existing approaches in Neo4j and RedisGraph, respectively.
+Among the tools, excluding Dinkel, GRev covers the most code (i.e., 597k lines) in Neo4j.
+Compared to GRev, Dinkel covers 25% more code.
+Benefiting from the powerful query generation, Dinkel can cover much deeper logic of GDBMSs that is related to processing advanced Cypher features and complicated data dependencies.
+In contrast, simple queries generated by existing tools rarely trigger such logic.
+In RedisGraph, excluding Dinkel, GDBMeter covers the most code (i.e., 19k lines).
+Compared to it, Dinkel covers 5% more code.
 The coverage improvement is less significant than in Neo4j, because RedisGraph supports fewer Cypher features and cannot handle some complicated Cypher semantics.
@@
 As a result, the space for Dinkel to improve the code coverage in RedisGraph is limited.
-Figure 17 shows the trends of the covered code in each GDBMS during testing.
-Dinkel and all the state-of-the-art cover new codes quickly and saturate in earlier stages.
-In almost all testing campaigns, Dinkel can cover more code than the existing tools.
 Found Bugs.
-As shown in Table 11, compared to existing approaches, Dinkel finds more bugs in both Neo4j and RedisGraph.
-We describe the relations of the found bugs in Figure 18.
-Note that none of these approaches have found logic bugs within 48 hours, so the bugs shown are related to internal errors or crashes.
-Specifically, Dinkel found 11 bugs in Neo4j.
-GDBMeter, GRev, and GAMERA found 1 bug, and this bug can also be found by Dinkel.
-In RedisGraph, Dinkel found 17 bugs, while GRev and GDBMeter found only 1 bug, and this bug can also be found by Dinkel.
-For the 27 bugs missed by existing approaches, their bug-triggering queries contain either complicated data dependencies or advanced Cypher features that are not supported by existing approaches.
-The bug shown in Figure 16 is one of the 27 bugs.
-These results demonstrate that Dinkel can find more bugs than existing approaches by generating more complex Cypher queries.
-Figure 18. Relation of bugs found by existing tools.
+Table 13 shows the number of bugs found by each approach over 5 runs.
+Figure 12 describes the bug relation.
+Dinkel finds the most bugs in both Neo4j (19 bugs) and RedisGraph (26 bugs), where logic bugs comprise 4 of the bugs found in Neo4j and 7 in RedisGraph.
+In Neo4j, GDBMeter, GRev, and GAMERA found 1 bug.
+In RedisGraph, GRev and GDBMeter found 1 bug.
+Both of these bugs were also identified by Dinkel.
+For the 43 bugs missed by existing approaches, their bug-triggering queries contain either complicated data dependencies or advanced Cypher features that are not supported by existing approaches.
+The bugs shown in Figure 10 and Figure 11 are two of the 43 bugs.
+These results demonstrate that Dinkel can find more bugs than existing approaches by generating and validating complex queries.
+Figure 12. Relation of bugs found by existing approaches.
 6. Related Work
 GDBMS Testing.
-GDBMS testing is an emerging research field and several approaches (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zheng et al., 2022; Zhuang et al., 2023; Mang et al., 2024) are proposed.
-All these approaches integrate techniques for constructing test oracles to identify bugs, especially logic bugs (i.e., the bugs cause GDBMSs to return incorrect results).
+GDBMS testing is an emerging research field, where several approaches (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zheng et al., 2022; Zhuang et al., 2023; Mang et al., 2024b) have been proposed.
+All these approaches aim to construct test oracles to identify logic bugs.
 Grand (Zheng et al., 2022) targets GDBMSs using the Gremlin query language and detects logic bugs using differential testing (McKeeman, 1998).
-Several approaches (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zhuang et al., 2023; Mang et al., 2024) support GDBMSs using Cypher.
+Several approaches (Jiang et al., 2023b; Kamm et al., 2023; Hua et al., 2023; Zhuang et al., 2023; Mang et al., 2024b) support GDBMSs using Cypher.
 Similar to Grand, GDsmith (Hua et al., 2023) also uses differential testing to find logic bugs.
-GDBMeter (Kamm et al., 2023) leverages the idea of query partitioning (Rigger and Su, 2020b), which decomposes the predicate of a query and checks whether the queries with the decomposed predicates produce consistent results with the original query.
-GraphGenie (Jiang et al., 2023b) modifies the graph patterns used in a query and checks whether the query with modified graph patterns satisfies the expected relationship (i.e., equivalence, subset, or superset) with the original queries.
-GAMERA (Zhuang et al., 2023) extracts metamorphic relations (Chen et al., 1998) from the manipulated graph data and validates GDBMSs by checking whether the outputs of generated queries satisfy these relations.
-Different from all these existing approaches, which focus on new test oracles for finding logic bugs, Dinkel focuses on improving query generation, which is the groundwork of GDBMS testing.
-An interesting future work is to combine the techniques of Dinkel and existing approaches, enhancing both query generation and test-oracle construction for finding more bugs in GDBMSs.
+GDBMeter (Kamm et al., 2023) leverages query partitioning (Rigger and Su, 2020b), which decomposes the predicate of a query and checks if the queries with the decomposed predicates produce consistent results with the original query.
+GraphGenie (Jiang et al., 2023b) modifies the graph patterns used in a query and checks if the query with modified graph patterns satisfies the expected relationship (i.e., equivalence, subset, or superset) with the original queries.
+GAMERA (Zhuang et al., 2023) extracts metamorphic relations (Chen et al., 1998) from the manipulated graph data and checks whether the outputs of generated queries satisfy these relations.
+Dinkel is different from all existing approaches in two key aspects.
+First, Dinkel significantly improves query generation, the core foundation of GDBMS testing.
+Second, while existing approaches support correctness testing for only queries following specific patterns, Dinkel generally tackles the problems of test-oracle construction, enabling validation for arbitrary queries.
 RDBMS Testing.
 Compared to GDBMS testing, testing relational database management systems (RDBMSs) is more mature, where both query generation (SQLsmith: A random SQL query generator., 2023; Zhong et al., 2020; Fu et al., 2022; Jiang et al., 2023a; Liang et al., 2023) and test-oracle construction (Rigger and Su, 2020c, a, b; Hao et al., 2023; Song et al., 2023; Tang et al., 2023; Jiang and Su, 2024) are well-researched.
-We focus on discussing the work on query generation for RDBMSs.
-SQLsmith (SQLsmith: A random SQL query generator., 2023) embeds the SQL grammar (SQL standard., 1992) and can generate queries containing complex statements.
-SQUIRREL (Zhong et al., 2020) integrates a new intermediate representation (IR) for modeling the dependencies among statements in a SQL query.
-Therefore, SQUIRREL can generate queries containing multiple statements.
-DynSQL (Jiang et al., 2023a) incrementally generates SQL statements for a query.
-For each statement, DynSQL queries the tested RDBMS to capture the latest DBMS state.
-In this way, DynSQL can generate complex and valid queries with multiple statements.
-LEGO (Liang et al., 2023) proposes type-affinity to describe the pattern of composing two types of SQL statements.
-Based on the type-affinity extracted from existing queries, LEGO can mutate and synthesize new queries that have a higher chance of increasing the code coverage of RDBMSs.
-Different from these approaches, Dinkel is designed for GDBMSs and integrates techniques for improving Cypher query generation.
+SQLsmith (SQLsmith: A random SQL query generator., 2023) embeds the SQL grammar (SQL standard., 1992) and can generate complex SQL statements.
+DynSQL (Jiang et al., 2023a) incrementally generates SQL statements for a query by querying the latest DBMS schema, which enables DynSQL to decouple the generation of each statement and helps it generate queries containing multiple complex statements.
+To find logic bugs in RDBMSs, PQS (Rigger and Su, 2020c) synthesizes customized queries that fetch specific rows of tables.
+If the tested RDBMS fails to fetch the rows, a logic bug is identified.
+Pinolo (Hao et al., 2023) modifies the predicate of a query, where the modified query should produce the subset/superset of the results of the original query.
+Otherwise, Pinolo identifies a logic bug.
+EET (Jiang and Su, 2024) transforms the expressions of queries in a semantic-preserving manner and checks whether the queries with the transformed expressions produce the same results as the original queries.
+EET identifies a bug if their results are not the same.
+Due to the separation of querying and data manipulation in the SQL language, EET cannot perform clause-level transformations the same way Dinkel can for Cypher queries.
+Unlike these approaches for SQL queries, Dinkel is designed for validating Cypher queries.
+Leveraging the features of the Cypher query language, Dinkel integrates novel techniques for both query generation and test-oracle construction.
 State-Aware Fuzzing.
-Fuzzing is a promising technique for detecting bugs in various software (Jiang et al., 2020, 2022; American Fuzzy Lop., 2021; Lyu et al., 2019; Yun et al., 2018; Padhye et al., 2019; Chen and Chen, 2018).
+Fuzzing is a promising technique for finding bugs in software (Jiang et al., 2020, 2022; American Fuzzy Lop., 2021; Lyu et al., 2019; Yun et al., 2018; Chen and Chen, 2018).
 Some approaches (Ma et al., 2023; Kim et al., 2022; Jiang et al., 2023a; Zhao et al., 2022; Atlidakis et al., 2019) have been proposed to find bugs more efficiently in state-sensitive systems.
-RESTler (Atlidakis et al., 2019) analyzes the API specification of the tested cloud service and generates request sequences that follow the inferred producer-consumer dependencies.
-RESTler also collects the response observed during prior executed requests to guide subsequent request sequence generation.
+RESTler (Atlidakis et al., 2019) analyzes API specifications of the tested cloud services and generates request sequences that follow the inferred producer-consumer dependencies.
+RESTler also collects the response observed during prior requests to guide subsequent request generation.
 LOKI (Ma et al., 2023) proposed to test blockchain consensus protocols.
@@
 StateFuzz (Zhao et al., 2022) is designed to test Linux drivers.
-It utilizes static analysis to recognize critical variables that affect the control flows or memory accesses, and represents program states using these variables.
-StateFuzz prioritizes test cases triggering new states using three-dimensional feedback.
-To effectively test USB gadget stacks, FuzzUSB (Kim et al., 2022) extracts the internal state machines from USB gadget drivers via static analysis and symbolic execution, and uses such state information as fuzzing feedback to guide test-case generation.
-Similar to these approaches, Dinkel also leverages state information and is the first state-aware testing approach designed for detecting bugs in GDBMSs.
+It utilizes static analysis to recognize critical variables that affect control flows or memory accesses, and represents program states using these variables.
+StateFuzz prioritizes test cases that trigger new states.
+For effectively testing USB gadget stacks, FuzzUSB (Kim et al., 2022) extracts the internal state machines from USB gadget drivers via static analysis and symbolic execution, before using this state information as fuzzing feedback for guiding test-case generation.
+Different from these approaches, Dinkel models GDBMS-related state information for allowing query generation to introduce complex data dependencies, all while retaining query validity.
 7. Conclusion
-In this paper, we have presented a novel, practical approach for testing GDBMSs using complex Cypher queries.
-Our approach models the graph state into two categories: query context and graph schema.
-Based on the model, we have proposed state-aware query generation, which incrementally constructs the clauses for the generated queries and intermittently updates the state information after each clause is constructed.
-We have implemented our approach as an automatic GDBMS testing framework, Dinkel, and evaluated it on three popular GDBMSs.
-The evaluation results demonstrate that Dinkel can effectively generate complex Cypher queries with high validity.
-Dinkel found 60 bugs, many of which had been missed by existing approaches.
-Considering its effectiveness in Cypher query generation, we believe Dinkel can be the foundation for GDBMS testing, facilitating and inspiring follow-up research, like Csmith (Yang et al., 2011) for compiler testing.
+We have presented a novel and practical framework, Dinkel, for validating GDBMSs.
+We model the graph state as query context and graph schema, and propose state-aware query generation to generate complex and valid Cypher queries for testing the deep logic of GDBMSs.
+Moreover, we propose two fine-grained query transformations: clause-level transformations and expression-level transformations, which can operate on arbitrary Cypher queries to validate their correctness.
+In our evaluation, Dinkel found 127 bugs in three well-known GDBMSs, among which 33 are logic bugs.
+Considering its significant advancement in bug finding, we believe Dinkel can lay a practical foundation for GDBMS testing, facilitating and inspiring follow-up research on GDBMS reliability.